### v2

# To CoT or not to CoT? Chain-of-thought helps mainly on math and symbolic reasoning

Abstract
Chain-of-thought (CoT) via prompting is the de facto method for eliciting reasoning capabilities from large language models (LLMs). But for what kinds of tasks is this extra “thinking” really helpful? To analyze this, we conducted a quantitative meta-analysis covering over 100 papers using CoT and ran our own evaluations of 20 datasets across 14 models. Our results show that CoT gives strong performance benefits primarily on tasks involving math or logic, with much smaller gains on other types of tasks. On MMLU, directly generating the answer without CoT leads to almost identical accuracy as CoT unless the question or model’s response contains an equals sign, indicating symbolic operations and reasoning. Following this finding, we analyze the behavior of CoT on these problems by separating planning and execution and comparing against tool-augmented LLMs. Much of CoT’s gain comes from improving symbolic execution, but it underperforms relative to using a symbolic solver. Our results indicate that CoT can be applied selectively, maintaining performance while saving inference costs. Furthermore, they suggest a need to move beyond prompt-based CoT to new paradigms that better leverage intermediate computation across the whole range of LLM applications [^1] Our code can be found at https://github.com/Zayne-sprague/To-CoT-or-not-to-CoT..
Figure 1: Left: meta-analysis of CoT literature; each point is a reported delta of CoT over direct answering for some (LLM, task) pair. Right: average performance of using zero-shot CoT v.s. direct answer prompts across five general reasoning categories, covering 20 datasets with 14 LLMs evaluated on each. In both sets of results, math and other kinds of symbolic reasoning are the domains that consistently see substantial improvements from CoT (red dotted line indicates the mean improvement from CoT across experiments).
1 Introduction
Chain-of-thought (CoT) (Nye et al., 2022; Wei et al., 2022) has become a widely used prompting technique for eliciting reasoning from language models. CoT can provide human-readable explanations of how problems are solved (Joshi et al., 2023; Lanham et al., 2023), but most frequently it is invoked to improve an LLM’s ability to answer complex questions via intermediate computation (Madaan & Yazdanbakhsh, 2022; Wang et al., 2023a; Dziri et al., 2023). Current post-training schemes for LLMs heavily infuse CoT capabilities into models: systems like ChatGPT or Llama 3.1 default to CoT when given reasoning problems (OpenAI, 2023; Dubey et al., 2024).
CoT has seen widespread usage, but it is most heavily explored in the domain of mathematical reasoning (Zhou et al., 2023a; Fu et al., 2023; Chae et al., 2024; Xu et al., 2024b; Qi et al., 2024). In fact, many “reasoning” methods for LLMs are evaluated only in the math domain; for instance, Lightman et al. (2024) frame their paper as “complex multi-step reasoning” and Mixtral-Large2’s release  [^2] https://mistral.ai/news/mistral-large-2407/ cited effort “enhancing the model’s reasoning capabilities”, but performance is only reported on GSM8K and MATH. CoT is reported to be effective across a wide range of studies, but many of these studies focus on a narrow slice of the task space. In areas beyond math, results show that CoT is not as useful (Kambhampati et al., 2024a) or can even hurt performance (Wang et al., 2024).
In this work, we aim to evaluate where prompt-based CoT helps and why. We begin with a systematic meta-analysis of recent literature that reports performance of CoT versus direct answering (DA). We then augment this picture by conducting experiments on 20 datasets and 14 contemporary LLMs across zero-shot and few-shot prompt settings. Finding 1: CoT only helps substantially on problems requiring mathematical, logical, or algorithmic reasoning. Figure 1 shows this holds both across the literature and our own experiments. We find only a few cases of large gain in other kinds of tasks, and many of these outliers feature some component of symbolic reasoning. For instance, on MMLU (Hendrycks et al., 2021a) and MMLU Pro (Wang et al., 2024), we analyze the improvements from CoT and find that CoT only gives benefit on math slices of the dataset. As much as 95% of the total performance gain from CoT on MMLU is attributed to questions containing “=” in the question or generated output. For non-math questions, we find no features to indicate when CoT will help.
How can we better understand why CoT improves on these questions and only these questions? The math and formal logical reasoning datasets we consider can be broken down into two stages of processing: a planning step (e.g., parsing a problem into equations) and an execution step (building intermediate outputs and working towards a solution) (Ye et al., 2023; Wang et al., 2023b; Sun et al., 2024). Finding 2: CoT primarily helps with the execution step that performs computation and symbolic manipulation, but falls short of what LLMs with tool augmentation can do. We find that LMs prompted with CoT can generate executable formal solution plans and execute those plans better than direct answering. But using LMs to generate a solution plan and then using an external symbolic solver to solve the plan outperforms using CoT for both steps for these tasks.
These results paint a picture that CoT’s utility is often circumscribed by tool augmentation: on problems where CoT helps, we already have more powerful tools than CoT that we can employ, and on “soft reasoning” problems like commonsense where no tools exist, we see limited benefit from CoT. This characterization has two major implications. First, CoT is unnecessary for many problems where it is widely employed: there exist more efficient prompting strategies that yield similar performance for much lower inference cost. Second, we see a critical need to move beyond prompt-based CoT to more sophisticated approaches based on search, interacting agents, or models more heavily fine-tuned for CoT. Future work can explore how intermediate computation can be better used to solve challenging problems outside of the math and symbolic reasoning domains.
2 Background: Chain-of-thought
The tasks we consider in this work consist of a question $\mathbf{q}\in\Sigma^{*}$ for a vocabulary $\Sigma$ and an answer $a\in\mathcal{L}(\mathbf{q})$ for a label set $\mathcal{L}(\mathbf{q})$. $\mathcal{L}(\mathbf{q})$ can consist of a data type like boolean or integer, classification labels, or problem-dependent labels like names of entities from $\mathbf{q}$. One exception that we still explore is BiGGen Bench (Kim et al., 2024), which instead relies on an LLM-as-a-judge (Dubois et al., 2023; Zheng et al., 2024b) to provide a label for generated long-form responses.
Prompting and chain-of-thought for reasoning
A large language model places distributions over strings $p(\mathbf{y})=\prod_{i=1}^{n}p_{\mathrm{LM}}(y_{i})$ where $\mathbf{y}\in\Sigma^{*}$. In practice, we can interpret these as conditional distributions $p(\mathbf{y}\mid\mathbf{x})$ where $\mathbf{x}$ is a user’s prompt. Typical invocation of an LLM involves forming a prompt $\mathcal{I}(\mathbf{q})$ that wraps the question with additional instruction, then drawing a sample response $\tilde{\mathbf{y}}\sim p(\mathbf{y}\mid\mathcal{I}(\mathbf{q}))$, and finally returning $a=\mathrm{extract}(\tilde{\mathbf{y}})$ using some kind of answer extractor.
For the tasks we consider in this work, the output $\tilde{\mathbf{y}}$ can take one of two forms. A direct answer only contains a string realization of $a$; e.g., $\mathbf{y}=(\mathrm{\_185},\mathrm{4})$ which is detokenized as the answer $a=1854$. A chain of thought is a longer sequence $\mathbf{y}$ including other tokens beyond the answer, e.g., $\mathbf{y}=(\mathrm{\_185},\mathrm{6},\mathrm{\_minus},\mathrm{\_2},\mathrm{\_equals},\mathrm{\_185},\mathrm{4})$. In both cases, the $\mathrm{extract}$ function must parse and detokenize the output; in CoT, there is some extra work to spot where the answer is placed.
Our prompts can explicitly encourage use of direct answer or chain of thought as strategies, which we denote as $\mathcal{I}_{\mathrm{da}}$ and $\mathcal{I}_{\mathrm{cot}}$. For eliciting CoT, this includes strategies like telling a model to “think step by step(Kojima et al., 2022). For directly answering a question, a prompt may say “immediately generate the answer”. We track the average location of the answer in the generated output for both CoT and direct prompts in Appendix F.3 to ensure that direct answer prompts give the answer early in the output. We also ensure that $\mathrm{extract}$ can parse answers from the generated output for each model, prompt, and dataset combination used in our experiments, tailoring the $\mathrm{extract}$ function as needed to ensure low unparseable rates for each model and task. [^3] We exclude a number of other “CoT-like” approaches in our analysis such as decomposed prompting (Khot et al., 2023; Zheng et al., 2024a) and multi-agent debate (Du et al., 2023; Chen et al., 2024). We focus on single prompt approaches. We deal with tool-augmented approaches in Section 5. All prompts and outputs per dataset per model have been uploaded to Huggingface and we include examples of some of our prompts in the Appendix J. We also experiment with few-shot CoT prompts, which we find perform similarly to zero-shot prompts; details about these are given in Appendix E.
Symbolic reasoning
Of key importance to this work is whether problems feature symbolic reasoning or not. We consider a problem to be symbolic if it can be grounded in a natural, well agreed-upon formal system. “$12\times 4$” is an example of a symbolic problem, which can be grounded in mathematics. Other systems include first-order logic (Saparov & He, 2023; Hua et al., 2024) or planning languages (Liu et al., 2023a; Valmeekam et al., 2023). Formally, for symbolic problems, we define a function $f$ that acts as a map that produces some symbolic expression $\mathcal{S}=f(\mathbf{q})$ from the question. $\mathcal{S}$ can be used as input for a solver to derive an answer, $\hat{a}=\mathrm{solve}(\mathcal{S})$.
Conversely, a problem like where on a river can you hold a cup upright to catch water on a sunny day? from CommonsenseQA (Talmor et al., 2019) is non-symbolic by our definition. While this problem could be formalized with some kind of predicate logic (Zhou et al., 2022; Quan et al., 2024; Zhou et al., 2024) or grounded in some kind of physical simulation (Hao et al., 2023; Wong et al., 2023), there is not a natural nor well agreed-upon framework for solving it.
We view non-symbolic to symbolic reasoning as a spectrum. MuSR (Sprague et al., 2024) is a “semisymbolic” dataset in that it does contain an underlying formal system (e.g., for its murder mysteries portion, the notion that $\mathrm{motive}(X)\wedge\mathrm{means}(X)\wedge\mathrm{opportunity}(X)\implies\mathrm{murderer}(X)$), but also involves substantial commonsense reasoning that does not map onto a formal system. In these cases, we can still form $\mathcal{S}=f(\mathbf{q})$, but $f$ must rely heavily on a language model and instantiate new information for $\mathcal{S}$ that is not directly represented in $\mathbf{q}$.
Central claim
Figure 1 shows that there are a large number of positive results on CoT reported in the literature. Informally, we believe many readers of the literature to hold the following view: $\mathcal{I}_{\mathrm{cot}}$ will outperform $\mathcal{I}_{\mathrm{da}}$ on nearly all reasoning problems, whether those problems involve symbolic or non-symbolic reasoning. Our evidence does not support this conjecture. We will show that this performance boost is strongest for symbolic and semi-symbolic tasks, while giving little to no improvement (or even hurting performance) on non-symbolic tasks.
3 Results from the Literature
Criteria and Process
We investigate all papers from ICLR 2024, a representative ML venue, and two representative NLP venues, EACL 2024 and NAACL 2024 (including Findings and Workshop papers). This resulted in 4,642 papers total that filtered using automatic and manual methods to papers including experiments comparing chain-of-thought, $\mathcal{I}_{\mathrm{cot}}$, against direct answering prompts, $\mathcal{I}_{\mathrm{direct}}$. A total of 110 papers were found that matched our criteria with 1,218 experimental comparisons. We then grouped the comparisons by the types of tasks and datasets being evaluated. More details on our automatic and manual filtering, as well as our categorization, can be found in Appendix A and B.
Table 1: A few categories for experimental comparisons. Full list in Appendix B.
Category Description
Symbolic and algorithmic Tasks involving symbol manipulation which can be solved by executing a program. This includes entity tracking datasets (e.g., SCONE, Coin Flip) and algorithmic tasks (e.g., BBH word sorting or finding shortest paths in a graph).
Math Tasks requiring mathematical reasoning, from grade-school math to advanced mathematics, including physics questions.
Logical reasoning Tasks designed to test for logical reasoning, whether deductive (Saparov & He, 2023, PrOntoQA), inductive (Bowen et al., 2024) or analogical (Ma et al., 2024) reasoning, including syllogisms and logical puzzles.
Encyclopedic knowledge Tasks requiring expert-level in-depth knowledge beyond mere commonsense, usually in an open-book setting.
Mixed datasets Datasets containing a variety of tasks, such as BIG-Bench Hard (BBH) or MMLU.
Figure 2: Results from our meta-analysis (grey dots) aggregated by paper and category (blue dots).
Results
Figure 2 shows the distribution of CoT deltas (CoT prompt minus the direct answer prompt performance) across our categorization of different task types found in the literature. Compared to Figure 1, we take the mean results per paper per category, indicated by blue dots, showing the trend across papers in the literature. The categories are ranked in order of ascending median CoT delta. The three categories which benefited the most from CoT are symbolic reasoning, math, and logical reasoning, with average improvements of 14.2, 12.3, 6.9, respectively. Average performance on these top three tasks with CoT was 56.9, whereas performance without CoT was 45.5. For other categories, the average performance with CoT was 56.8, compared to 56.1 without CoT. We do not consider this small improvement a victory for CoT. CoT involves more computation than direct answering, and a truly fair comparison between the methods should match the compute of the two methods, e.g., ensembling across multiple prompts.
Do any non-math datasets benefit from CoT?
On the right side of Figure 2, we show the top 10 outliers from our observed trend, namely papers with high CoT deltas averaged across experiments in tasks other than math, symbolic, or logical reasoning. Although not categorized as math or logic, several of these are related to logical, mathematical or symbolic reasoning in some way. From this list, the dataset which benefits the most most from CoT is BIG-bench Hard (BBH) (Suzgun et al., 2023), a benchmark consisting largely of problems requiring algorithmic, arithmetic or logical reasoning. For instance, BIG-bench Navigate is a spatial reasoning task, but relies heavily on a mathematical primitive of counting steps taken to derive a final conclusion. Similarly, while BIG-bench Temporal is a temporal reasoning task (answering questions about when certain events could have occurred), it requires deductive reasoning to solve. In addition, Legal Argument Reasoning (SemEval-2024 Task 5) (Bongard et al., 2022) was categorized as context-aware QA, but also requires substantial reasoning ability. Finally, MMLU-Moral Scenarios (Hendrycks et al., 2021a) requires answering two independent questions at once, which essentially involves a symbolic combination of two simpler questions.
There are a few outliers that less clearly follow the trend. ScienceQA (Lu et al., 2022) consists of multiple choice questions across a range of natural and social science disciplines, though it is hard to interpret gains without knowing breaking down performance by subject or question type. The dialogue evaluation dataset from Jia et al. (2024) sees large improvements with CoT, but this is a proprietary dataset, and we note that other essay scoring results in our meta-analysis (Li et al., 2024; Stahl et al., 2024) did not show improvements with CoT. Other non-math, symbolic or logical datasets that benefit from CoT are Commitment Bank (de Marneffe et al., 2019) and the task of eliciting verbalized confidence (Xiong et al., 2024). Nevertheless, these are exceptions to the rule. The majority of the reported benefits from using CoT in the NLP and ML literature comes from math or math-related tasks.
4 Results from Experiments
4.1 Experimental Setup
Dataset, Models, Prompts
All datasets, models, and prompts we evaluate over can be found in detail in the tables 3, 4, and 5 of Appendix C. We restricted our experiments to English models commonly used and benchmarked on general reasoning datasets. Our datasets include those which are widely used in CoT and reasoning literature, including a mix of non-symbolic, semisymbolic, and symbolic reasoning. They span different formats, including multiple-choice, short-answer, and free-response; however, most of these datasets are multiple choice or short answer, as CoT is not typically used in long-form response settings. We also categorize each dataset into a larger category of reasoning required to solve it: Commonsense, Knowledge, Symbolic, Mathematical, and Soft Reasoning. We define Soft Reasoning as questions relying on commonsense and natural language but going beyond simple inferences about these statements. Finally, we explore several prompting strategies for eliciting reasoning from language models, as past work has emphasized the importance of the prompt (Yang et al., 2024). However, we generally found slight performance differences; see Appendix D for details. We therefore focus on prompts similar to Kojima et al. (2022) and Wei et al. (2022) for zero-shot and few-shot settings, respectively, with alterations to improve the model’s ability to produce desired behavior (i.e., formats that allow for easily parsed answers). We upload all our prompts and outputs for each model for each prompting strategy on Huggingface [^4] https://huggingface.co/collections/TAUR-Lab/cot-analysis-project-66bbb9e5e0156e65059895f5.
Implementation Details
We use a high-throughput inference package, vLLM (Kwon et al., 2023), for the model inference process. We use greedy decoding on all models. Our prompts are taken from the Llama 3.1 evaluations when available (Dubey et al., 2024), and minor adjustments are made to unify prompting strategies. For other datasets, we either use the standard prompt for the dataset from the corresponding original paper or implement our own prompt. Our answer parser ($\mathrm{extract}$) is tailored to each dataset and model. Specific details about each dataset, its prompts, and answer extractor can be found in Appendix C.
Figure 3: Left: Performance gain from using CoT for each reasoning category. Right: Performance gain from using CoT for each dataset, averaged across models and broken out across 5 representative models. Red lines indicate median improvement. In both plots we see a consistent trend: most improvements from using CoT are from math and symbolic reasoning.
4.2 Results
Where does zero-shot CoT improve over direct prompts?
On datasets that require math (MATH, GSM8K) or formal logic (ContextHub, MuSR to a lesser degree) to answer the problem.
Figure 3 on the left shows the average CoT performance improvement for each reasoning category from Figure 1 (right); raw numbers can be found in Table 6 of the Appendix. On the right, Figure 3 shows the performance gain from using CoT for each dataset, averaged across all models and for a selection of individual models. On non-symbolic reasoning categories and datasets, specifically those that contain questions primarily involving commonsense (CSQA, PIQA, SiQA), language understanding (WinoGrande), and reading comprehension (AGI LSAT, ARC-Easy, ARC-Challenge), there is little to no separation between the performance of zero-shot CoT and zero-shot direct answer. Despite these datasets involving reasoning, CoT does not yield improvement.
By contrast, the mathematical and symbolic categories get larger boosts in improvements alongside symbolic and many semi-symbolic datasets. MATH and GSM8K show gains as large as 41.6% and 66.9%, respectively. The semi-symbolic datasets like ContextHub and MuSR Murder Mysteries show moderate gains. These datasets require the application of logical rules to reach the answer, e.g., first-order logic parsed from simple natural language (ContextHub) or more complex commonsense statements (MuSR Murder Mysteries). All results are shown in the Appendix F.1 as well as a full list of numeric results for both CoT and direct answer prompting in Table LABEL:tab:raw_exp_0_zs_results. We also explored the few-shot setting and found it had little impact on when CoT will help; see Appendix E.
Does the answer format impact where CoT will help?
Not much. Free response capabilities may be hindered by pre-planning or reasoning about the correct response.
Many of the commonly-used datasets for problems other than math are multiple choice. We highlight here that CoT has similar performance to direct answer across models for two datasets that are not multiple-choice and contain varying levels of non-symbolic reasoning. First, MuSiQue (Trivedi et al., 2022) is a short-form QA task requiring multi-hop reasoning. We consider this a semi-symbolic dataset as the questions have an explicit multi-hop structure. Because answer spans in MuSiQue can be paraphrased in many different ways, we use GPT-4o to judge if two answer spans are equivalent. Despite being semi-symbolic, we see no overall improvement from CoT.
Second, BiGGen Bench (Kim et al., 2024) uses free-form responses as the answer to a question, and an LLM-as-a-judge is used to evaluate these responses on a scale of 1 to 5. The free-form nature of the answers blurs the lines between CoT and direct answer. However, we devised a CoT prompt for this setting where we ask the language model to generate a plan for the free-form response (the reasoning part), and then we ask it to generate the full response (the answer part) all in one generation. We then only give the response to the judge. We use GPT-4o mini as the judge with the prompt from Kim et al. (2024). We also exclude slices from BiGGen Bench that ask the LLM to “Think step by step” within the question, as comparing it to direct answer is difficult with these prompts. We plot the performance of BiGGen Bench as the number of times a prompt receives a score of 4 or better on each question. CoT leads to mild overall improvement here, which we expect to see given the benchmark’s inclusion of reasoning questions (including several categories of math) and other categories such as planning.
Are the gains in Knowledge, Soft Reasoning, and Commonsense significant?
Mostly no, except for MMLU, StrategyQA, and MuSR.
We tested the significance of the improvements from CoT on the 13 datasets in the Knowledge, Soft Reasoning, and Commonsense reasoning categories using paired bootstrapping to assess whether CoT gives a significant improvement. To account for multiple comparisons, we applied a Bonferroni correction, setting the p-value to $0.00027$ to account for the 14 models and 13 datasets. About 32% (59) of the datasets that show a benefit in these three reasoning categories were considered significant. Nearly half of these comparisons (26) are on MMLU and MMLU Pro. On these datasets, we find that CoT is mainly helping on math-related questions. StrategyQA and MuSR also received a consistent performance boost across 10 and 6 models respectively. StrategyQA is often used to benchmark reasoning methods and is built specifically to get a benefit from methods that decompose the question into steps, so a gain in performance is not unprecedented. MuSR, similarly, was built to have multiple steps of complex natural language reasoning, which may receive benefits from CoT. The remaining datasets that receive significant benefits are spread across the datasets and models.
Why do MMLU and MMLU Pro get a boost?
MMLU and MMLU Pro contain many different questions requiring different types of reasoning. We separated MMLU and MMLU Pro questions into two bins, those related to math and those not related to math, by checking if the questions text or generated response from the LLM includes an “=”. Figure 4 shows that a majority of the performance gain seen from MMLU and MMLU Pro is from the math slices of each dataset. See more details in Appendix G.
Figure 4: CoT deltas between MMLU and MMLU Pro performance when a question or generated response contains an “=” (With =) or not (Without =). We filter out any questions that do not result in a final answer (degeneration, etc.). CoT primarily helps on the pairs of questions and generations that contain an “=”, which indicates math-related questions.
5 Strengths and weaknesses of CoT at formal reasoning
Previous sections establish that CoT primarily helps with symbolic reasoning tasks, but not why. Many symbolic and semi-symbolic tasks be broken down into two stages (Ye et al., 2023; Pan et al., 2023; Jiang et al., 2024): planning, either via a formal or informal specification via prompting (Sun et al., 2024; Wang et al., 2023b), and execution, using the same LM or external solvers. In this section, we attribute the performance gains from CoT on symbolic tasks to these two stages.
Given a question that requires symbolic reasoning, we define the planning stage as extracting all variables from the context into a formal specification and defining their relations. The execution stage uses a solver that takes as input a plan and can be run in an orderly fashion to derive the final answer. Using our notation from Section 2, let $f(\mathbf{q})=\mathcal{I}^{m}_{\mathrm{planning}}(\mathbf{q})$ be a mapping of the question $\mathbf{q}$ to a symbolic plan $\mathcal{S}_{\mathrm{plan}}$ that can be executed by the language model or by an external symbolic solver, $\hat{a}=\mathrm{solve}(\mathcal{S}_{\mathrm{plan}})$, where $\hat{a}$ is the final answer for $\mathbf{q}$.
By separating planning and execution in this way, we can test how much a language model can gain from only having a plan, to having a plan and solving it with CoT, or to having a plan and then solving it with an external symbolic solver. Given a plan $\mathcal{S}_{\mathrm{plan}}\sim\mathcal{I}^{m}_{\mathrm{planning}}(\mathbf{q})$, we compare the performance of the settings below to evaluate at which stage LM is most effective and falls short.
Figure 5: Prompt variants that separate planning and execution for GSM8K. For all prompt variants besides direct answer and CoT (not shown), we few-shot prompt an LLM to first generate a Python program as a solution plan. For Plan + Direct Solver, the LLM is prompted to directly give an answer from the plan; for Plan + CoT Solver, the LLM is prompted to solve the plan step-by-step with CoT and give an answer; for Plan + Tool Solver, we feed the plan into a Python interpreter.
5.1 Settings Evaluated
Settings 1 and 2: Few-shot direct answer and CoT: We use the few-shot direct answer and CoT prompts from Section 4.1 as baselines. Figure 5 includes an example of each setting on GSM8K.
Settings 3 and 4: Plan + Direct Solver and Plan + CoT Solver: Here we use inspiration from Xu et al. (2024a) and generate a symbolic plan using the same strategy as Ye et al. (2023). Specifically, we use a few-shot prompt $\mathcal{I}^{m}_{\mathrm{planning}}$ to generate a formal specification $\mathcal{S}_{\mathrm{plan}}$ that should be executable by a symbolic solver. In the same prompt LMs are asked to solve their generated specification $\mathcal{S}_{\mathrm{plan}}$ and derive the final answer $\tilde{\mathbf{y}}\sim p(\mathbf{y}\mid\mathcal{I}_{\mathrm{da}}(\mathcal{S}_{\mathrm{plan}}))$, either directly giving the answer after generating the specification (Plan + Direct Solver) or providing step-by-step explanations and tracking of intermediate steps for the derivation (Plan + CoT Solver). Particularly, $\mathcal{S}_{\mathrm{plan}}$ is a Python program for math datasets, and is a set of first-order logic specifications for logical reasoning datasets.
Setting 5: Plan + Tool Solver: We then evaluate how effective CoT can be at performing symbolic computations compared with external symbolic solvers. Following prior work on augmenting LMs with tools for math and logic questions (Ye et al., 2023; Pan et al., 2023; Gao et al., 2023; Chen et al., 2023), we generate $\mathcal{S}_{\mathrm{plan}}$ the same way as in CoT Solver, but now feed in the plan into a symbolic solver (Python interpreter or a SMT Solver), such that $\hat{a}=\mathrm{solve}(\mathcal{S}_{\mathrm{plan}})$.
Figure 6: Performance of prompt variants that separate planning and execution for math and logical reasoning datasets. Despite outperforming direct answer for solving a formal plan and deriving the final answer, CoT is still limited in performing symbolic computations: there is a large performance boost from Plan + Tool Solver over CoT and Plan + CoT Solver on average across all models.
Evaluation Setup: We compare the performance of each setting on math (GSM8K) and logical reasoning (ContextHub and FOLIO) datasets. We follow Gao et al. (2023) to include GSM8K-Hard, a minimally modified version that replaces numbers of GSM8K with larger numbers, to account for the possibility of recent LLMs overfitting GSM8K by data contamination (Zhang et al., 2024).
For Plan + Direct solver and Plan + CoT solver, we use the few-shot prompts from Ye et al. (2023). For Plan + Tool solver, we use state-of-the-art tool-augmented prompting methods. Particularly, for GSM8K, we use Program-aided Language Model (Gao et al., 2023, PAL) that executes the LM-generated plan with a Python interpreter. For logical reasoning datasets, we use Satisfiability-Aided Language Model (Ye et al., 2023, SatLM) that uses automated theorem prover Z3 (De Moura & Bjørner, 2008) to solve the generated specifications. If the generated plan cannot be parsed by the tool, we use random guessing when the question is multiple choice, and mark it incorrect otherwise.
5.2 Evaluation Results
Figure 6 shows the results across a representative selection of models. Detailed numerical results, including the unparseable rates of model-generated plans, can be found in Appendix H.
When comparing direct answer with Plan + Direct solver and Plan + CoT solver, we note that for many datasets and models, only having a plan does not account for most of the performance gain. Compared with direct answer, CoT or Plan + CoT solver is needed for strong performance. Tracking the execution with one of these methods gives the strongest accuracy benefit, especially for math-heavy datasets.
Despite their strength over direct answer and Plan + Direct solver, CoT and Plan + CoT solver are dominated by Plan + Tool solver in most settings. LLMs are limited by their ability to execute and track steps compared with symbolic solvers.
We argue that these results provide an explanation of why CoT helps on symbolic tasks. While all tasks could feasibly benefit from a detailed description of how to solve each individual question (e.g., a plan in the context of this section), CoT only outperforms direct answer when these steps require a substantial amount of tracing and computation. In these settings, we can see clear performance benefit from using symbolic solvers; CoT appears to be a poor (but universal) approximation to such solvers. When possible, LLMs should be paired with symbolic solvers at inference time when solving symbolic tasks to achieve consistently better performance over direct answer and CoT.
6 Discussion and Related Work
Where is CoT helping and why?
Our results showing CoT improvement for math and logic aligns well with early work on CoT for LLMs such as Scratchpads (Nye et al., 2022). As CoT gained popularity, its application has broadened to tasks that canonically do not require multiple steps. It can often yield small improvements over direct answering. We believe this led to the current prevailing sentiment that deliberation should improve performance on any task requiring some type of reasoning (our original claim from Section 2). However, our results show a clear separation between performance on non-symbolic and symbolic tasks. If, in theory, any question could benefit from deliberation, why is CoT only benefiting the questions that can be solved through symbolic manipulation? Our results from Section 5 suggest that the primary benefit of CoT comes in the ability to execute symbolic steps and track their output. Not all tasks have this feature: for example, questions from CommonsenseQA can hardly be translated into formally grounded and executable solution plans. Datasets like StrategyQA may feature multiple steps of reasoning, but executing those steps is not complex, so the benefits of CoT are small. It is unclear whether explicitly instilling models with particular modes of deliberation, like process of elimination for multiple choice questions, might make them more effective for non-symbolic tasks, or whether there’s a fundamental limitation imposed by their pre-training data. We leave this distinction for future work.
Can we improve CoT further?
Our work treats chain-of-thought variants that explicitly don’t involve multiple inferences. There is evidence that using additional calls to LLMs can help (Du et al., 2023; Yao et al., 2023; Besta et al., 2023; Chen et al., 2024), but these methods use significantly increased computation, and careful benchmarking sometimes reveals that naive techniques are as good as iterative ones (Olausson et al., 2024). However, past theoretical results show that Transformers are augmented in a fundamental way by CoT (Liu et al., 2023b; Merrill & Sabharwal, 2024); we believe this indicates the potential for improving CoT beyond prompt-based CoT. On the other hand, recent methods showing benefit from “internalizing” CoT (Deng et al., 2024) may indicate that explicit generation of intermediate tokens is not used to its full potential.
Limitations
One set of tasks we do not cover in our experiments (except for BiGGen Bench) is long-horizon planning. However, many works in the literature have already discussed the efficacy of planning with CoT. We also do not address the data contamination of some of these models on the datasets. We try to mitigate this by including multiple models, datasets (new and old), and our meta-analysis. For more discussion of planning and dataset contamination, see Appendix I.
7 Conclusion
In this work, we characterize the performance of prompt-based CoT through a meta-analysis of the literature and experiments across different models, datasets, and prompts. We find that CoT predominantly helps on math and formal logic, largely due to its ability to trace the intermediate steps of a problem. But CoT rarely outperforms tool-augmented approaches for these same problems. We believe that CoT remains a powerful technique, but to give improvement across a wider range of NLP tasks, research should move beyond prompt-based CoT to new paradigms like search, interacting agents, or better fine-tuned models.
Reproducibility
For our experiments, we provide in-depth details of how we evaluated models on each dataset in Section 4.1 and Appendix C. Furthermore, we release all prompts for every dataset on Huggingface, including per model output and sampling parameters. For our meta-analysis of the literature, we describe our filtering criteria and process of annotating experiments into high-level categories in Section 3 and Appendix B. We also release the full list of papers in our meta-analysis together with extracted experimental comparisons and task category annotations.
Appendix A Meta-analysis Expanded Details on Criteria and Process
Automatic Selection and Paper Filtering
We investigate all papers from ICLR 2024, a representative ML venue, and two representative NLP venues, EACL 2024 and NAACL 2024 (including Findings and Workshop papers). We filtered all 4,642 papers (2,259 from ICLR 2024 and 2,382 from the two ACL-affiliated conferences) for those with at least two occurrences of “CoT”, “chain-of-thought”, or “chain of thought”, resulting in 516 papers. There are conceivably papers using CoT called by another name (e.g., Scratchpads), but we believe these 516 give a representative sample appropriate for systematic analysis.
Manual Paper Filtering and Results Extraction
We then filter down to papers that perform a comparison of CoT prompting vs. direct prompting, whether or not this is core to the paper’s research question. We manually filtered the 516 papers in question and extracted the key results from those that remained. We excluded multimodal models, CoT-fine-tuned models, any experiments where the “CoT” method involves multiple forward passes (e.g., self-consistency (Wang et al., 2023c) and tree-of-thought (Yao et al., 2023)), [^5] These systems use more compute than direct answer, and there is not a clear comparison to be made here. Moreover, our anecdotal coverage of these methods shows that they are most used for math, coding, and logic settings, for which we already have high representation among reported CoT methods. and systems that augment LLMs with external tools (discussed more in Section 5).
For each paper passing through these criteria, we manually extracted the results from key tables comparing CoT and direct answer prompts. We only include results where the CoT and direct prompts are run on the same model and same dataset while being on a scale of 0 to 100 (excluding Likert scale evaluations, for example) for a more direct comparison. When papers include various CoT or direct answer prompts (including zero/few-shot variants), we always take the best-performing prompt for both. We focus on key test results where applicable, excluding dev sets if they are reported alongside test and also excluding numbers from ablations or nonstandard subsets of datasets.
This resulted in a total of 1,218 experimental comparisons across 110 papers (35 from ICLR and 75 from NAACL and EACL) covering 264 datasets. Details and more information will be available in our GitHub Repo.
Categorization
Given the large number of tasks and datasets being compared, we grouped each task into a set of 14 categories. These categories were determined based on the description (and possibly examples) of the task, not taking into account system performance. These categories abstract over traditional NLP task classifications (e.g., NER, reading comprehension) and take into account both the task format and the kinds of reasoning involved. Definitions for several categories are shown in Table 1 and the full description is given in Appendix B.
Appendix B Quantitative Meta-Analysis
See the full list of categories and their descriptions that we used for the meta-analysis in Table 2.
Category Description
Symbolic and algorithmic Tasks involving symbol manipulation which can be solved by executing a program. This includes entity tracking datasets (e.g., SCONE, Coin Flip) and algorithmic tasks (e.g., BBH word sorting or finding shortest paths in a graph).
Math Tasks requiring mathematical reasoning, from grade-school math to advanced mathematics, including physics questions.
Logical reasoning Tasks designed to test for logical reasoning, whether deductive (Saparov & He, 2023, PrOntoQA), inductive (Bowen et al., 2024) or analogical (Ma et al., 2024) reasoning, including syllogisms and logical puzzles.
Commonsense reasoning Datasets designed to test for commonsense knowledge and reasoning, i.e., world knowledge that most people would have, rather than specialized expert-level knowledge in a discipline acquired after years of study.
Encyclopedic knowledge Tasks requiring expert-level in-depth knowledge beyond mere commonsense, usually in an open-book setting.
Spatial and temporal reasoning Datasets designed to test for an understanding of space and spatial relations (e.g., navigation) or reasoning involving time and sequences over time.
Multi-hop QA Questions involving the composition of multiple steps of reasoning in order to arrive at an answer, such as “What is the capital of the country whose scientist discovered penicillin?”
Context-aware QA Tasks such as closed-book QA and reading comprehension involving reasoning about a given text in context. The context is often a short passage, but could also take the form of a knowledge graph (KBQA) or a table. This category also includes information extraction tasks, such as NER or relation extraction.
Entailment Tasks involving establishing the inferential relation between two texts, prototypically NLI, but also including fact verification.
Text classification Tasks involving the classification of a text into a small set of categories, such a topic or sentiment classification, but also involving tasks such as hate speech detection and misinformation detection.
Generation Tasks involving text generation, including machine translation, dialogue, question generation, as well as code generation. Tasks such as SQL execution (Lei et al., 2024) or systematic transformations of data (e.g., SCAN (Lake & Baroni, 2018)) are excluded because they can be solved by executing a program.
Meta-linguistic Tasks probing for models’ knowledge of linguistics, such as identifying the main subject of a sentence or solving linguistic puzzles.
Mixed datasets Datasets containing a variety of tasks, such as BIG-Bench Hard (BBH) or MMLU.
Other Tasks which did not fit in any of the other categories, such as evaluating AI safety, eliciting models’ verbalized confidence, or melody retrieval.
Table 2: Categories and their descriptions for the meta-analysis.
Appendix C Expanded Experimental Details
Models Llama 2 7B Chat (Touvron et al., 2023), Mistral 7B Instruct v0.3 (Jiang et al., 2023), Llama 3.1 8B Instruct (Dubey et al., 2024), Llama 3.1 70B Instruct, Gemma 2 9B It (Riviere & et. al, 2024), Phi-3 Small 8k Instruct (Abdin et al., 2024), gpt-4o-mini-2024-07-18*, gpt-4o-2024-08-06*, Gemini 1.5 Flash*(Reid & et. al, 2024), Gemini 1.5 Pro*(Reid & et. al, 2024), claude-3-haiku-20240307* (Anthropic, a), claude-3-5-sonnet-20240620* (Anthropic, b)
Datasets CommonsenseQA (Talmor et al., 2019), StrategyQA (Geva et al., 2021), SiQA Sap et al. (2019), PiQA (Bisk et al., 2019), Winogrande (Sakaguchi et al., 2021), GPQA (Rein et al., 2023), MuSR (Sprague et al., 2024), ContextHub (Levels 1 and 2 only) (Hua et al., 2024), ARC (Clark et al., 2018), AGIEval LSAT (Zhong et al., 2023), MMLU (Hendrycks et al., 2021a), MMLU Pro (Wang et al., 2024), MATH (Hendrycks et al., 2021b), GSM8K (Cobbe et al., 2021), GSM8K-hard (Gao et al., 2023), FOLIO (Han et al., 2022), MuSiQue (Trivedi et al., 2022), Big-Bench Hard (Suzgun et al., 2023; Srivastava et al., 2022), BiGGen Bench (Kim et al., 2024)
Prompts zero-shot direct answer, zero-shot CoT (Kojima et al., 2022), few-shot direct answer (Brown et al., 2020), few-shot CoT (Wei et al., 2022)
Table 3: Models, datasets, and prompting strategies used in our experiments. Models marked with † are run with a 4k context size window. Note that Gemma has a larger than 4k context size window, but VLLM only supports up to a 4k context size window for it. Models marked with * indicate closed-source models that cannot handle prefixed assistant messages. Datasets marked with $\triangle$ do not have a few-shot setting.
Dataset Type Answer Format $m$-Shots
CommonsenseQA Commonsense Multiple choice 7
StrategyQA Commonsense True or False 6
SIQA Commonsense Multiple choice 0
PIQA Commonsense Multiple choice 0
Winogrande Commonsense Multiple choice 0
Arc Easy Knowledge Multiple choice 0
Arc Challenge Knowledge Multiple choice 0
AGIEval LSAT Soft Reasoning Multiple choice 3
BiGGen-Bench Soft Reasoning Free response 0
MMLU Knowledge Multiple Choice 5
MMLU Pro Knowledge Multiple Choice 5
BigBench-Hard Symbolic Multiple Choice 0
MuSR Soft Reasoning Multiple Choice 1
GPQA Mathematical Multiple Choice 3
MuSiQue Soft Reasoning Short Answer 0
GSM8K Mathematical Short Answer 8
GSM8K-Hard Mathematical Short Answer 8
FOLIO Symbolic True, False, or Unknown 4
ContextHub Symbolic True, False, or Neither 3
MATH Mathematical Short Answer 4
Table 4: List of datasets used in our experiments. We categorize each dataset into one of five categories based on the type of reasoning required: Commonsense, Knowledge, Soft Reasoning, Symbolic, or Mathematical. We also report answer formats. When we use few-shot prompts, we mark how many examples those prompts contain. BiGGen Bench has many categories of questions that explicitly ask for CoTs in the response; we ignore those categories for our evaluation.
Model Context Length Is Open Source
Llama 2 7B Chat 4k True
Mistral 7B Instruct v0.3 8k True
Llama 3.1 8B Instruct 128k True
Llama 3.1 70B Instruct 128k True
Gemma 2 9B It 4k True
Qwen 7B Instruct 131k True
Qwen 72B Instruct 131k True
GPT4o-Mini 128k False
GPT4o 128k False
Gemini 1.5 Pro 128k False
Gemini Flash 1m False
Claude 3.5 Sonnet 200k False
Claude 3 Haiku 200k False
Table 5: List of models for our experiments. We focus on contemporary instruction-tuned models; although pretrained and smaller language models could be used, they are not the focus of our study. Prompts and outputs used for each model are available on Huggingface. Note that Gemma can accept more than 4k input tokens, but we are restricted to 4k by vLLM.
A full list of the datasets can be found in Table 4. Each model can be seen in Table 5. We use one answer parser for all datasets of the same answer response format (one for multiple choice, short answer, etc.); however, some datasets require special handling and have edge cases that we handle separately from the rest of the datasets. Similarly, for each model, we use the exact same prompt across them, except when closed source models require different prompts because they do not allow for partial completions (i.e., when we cannot put “let’s think step by step” to warm-start the assistant’s response). All prompts are given in our Huggingface repo, including the model output and what our answer parser extracted as the answer.
Experiments were conducted either by invoking APIs or by running open-source models on our own hardware, mostly on a machine with 8 A40s or 4 Quadro RTX 8000s. All locally hosted models were hosted with vLLM. All parameters given to the vLLM API endpoint are given in the Huggingface repo as well.
Appendix D Other CoT prompt variants
D.1 Testing Performance Volatility Across Prompts
To test the impact of prompt choice on performance, we performed our zero-shot experiment on Llama 3.1 8B with 7 different datasets and 4 different zero-shot CoT prompting strategies common in the literature (Kojima et al., 2022; Wang et al., 2023b; Zhou et al., 2023b; Yang et al., 2024). Figure 7 shows variation due to prompts is typically small and no prompt gives a consistent gain over the other. For our experiments, this suggests that different prompts have small effects on the overall outcome on average.
Figure 7: Performance of multiple prompts commonly used to elicit reasoning through CoT in the zero shot setting. Each prompt starts the assistant completion with a different phrase meant to elicit reasoning. All results are from using Llama 3.1 8B Instruct. For the Kojima variant, we explicitly place “Let’s think step by step.” in the assistant message. There is very little variation between the CoT prompts on average.
Appendix E Few-shot Experiments
Compared to a zero-shot prompt, a few-shot prompt additionally contains demonstrations of the relevant reasoning mode on different problem instances $\{(v(\mathbf{q}_{i}),\mathbf{y}^{*}_{i})\}$. Few-shot prompts for direct answer simply encode the answer $a_{i}$ as $\mathbf{y}^{*}_{i}$, whereas few-shot prompts for chain-of-thought include a reasoning trace ending in the correct answer. Now we can define the $m$-shot direct prompt as $\mathcal{I}_{\mathrm{da}}^{m}(\mathbf{q})=v_{\mathrm{da}}(\mathbf{q}_{1})\mathbf{a}_{1}v_{\mathrm{da}}(\mathbf{q}_{2})\mathbf{a}_{2}\ldots v_{\mathrm{da}}(\mathbf{q}_{m})\mathbf{a}_{m}v_{\mathrm{da}}(\mathbf{q})$ and the $m$-shot cot prompt as $\mathcal{I}_{\mathrm{cot}}^{m}(\mathbf{q})=v_{\mathrm{cot}}(\mathbf{q}_{1})\mathbf{y}^{*}_{1}v_{\mathrm{cot}}(\mathbf{q}_{2})\mathbf{y}^{*}_{2}\ldots v_{\mathrm{cot}}(\mathbf{q}_{m})\mathbf{y}^{*}_{m}v_{\mathrm{cot}}(\mathbf{q})$.
Figure 8: Average performance improvement from using CoT across different models in the zero-shot and few-shot settings. Each bar represents how much CoT improves the accuracy for that specific setting. In general, CoT in the few-shot setting does not change the qualitative performance of CoT versus zero-shot, though it can change the magnitude for symbolic datasets.
Figure 8 shows the difference between few-shot prompting and the zero-shot setting discussed in the main text of the paper. We see that using CoT in the few-shot setting largely does not change the datasets that benefit from it. Only one dataset, MuSR Team Allocation, starts to improve with few-shot; however, we believe this to be an exception because the final step to derive the answer is complex in the prompt and clearer in the examples. The magnitude of improvement over direct answer prompting when using CoT is also similar to the zero-shot setting.
Appendix F Expanded CoT vs Direct Experimental Results
F.1 Full Zero-shot Results
Model Commonsense Knowledge Mathematical Symbolic Soft
DA % CoT % DA % CoT % DA % CoT % DA % CoT % DA % CoT %
Claude-3 Haiku 74.3 77.2 73.0 76.1 18.1 48.2 38.6 48.7 55.9 56.6
Claude-3.5 Sonnet 84.3 85.8 83.8 88.8 38.7 59.0 53.2 67.1 67.6 75.7
GPT-4o Mini 81.8 83.2 73.6 83.1 22.9 59.7 48.1 60.9 61.1 63.5
Gemini 1.5 Flash 80.3 76.8 78.2 81.0 27.2 55.7 47.0 59.7 60.6 62.6
Gemini 1.5 Pro 80.4 78.3 80.9 83.8 35.4 58.5 52.9 62.6 64.1 67.8
Gemma 2 9b 75.0 76.1 74.9 76.9 18.5 50.5 46.7 55.8 58.2 60.5
Gpt-4o 87.3 87.7 82.9 88.6 36.5 63.3 55.7 68.3 65.9 74.0
Meta-Llama 2 7b 51.4 50.9 44.1 46.6 9.3 17.2 22.4 35.4 37.2 37.6
Meta-Llama 3.1 70b 84.2 84.7 82.4 85.6 24.9 54.9 49.0 60.0 65.7 69.5
Meta-Llama 3.1 8b 72.9 73.4 70.1 74.1 16.0 47.8 34.8 51.6 55.0 56.2
Mistral 7b 58.3 61.8 62.0 64.5 10.9 28.9 41.8 45.0 48.6 49.7
Phi-3 Small 8k 70.8 72.5 76.1 79.7 17.8 47.1 51.2 58.7 57.9 56.4
Qwen 2 72b 82.9 84.9 78.6 84.6 23.9 58.5 48.2 58.7 64.2 65.1
Qwen 2 7b 64.0 66.1 65.2 71.3 15.9 53.5 43.8 52.3 54.4 49.4
Average 74.8 75.7 73.3 77.5 22.6 50.2 45.2 56.1 58.3 60.3
Table 6: Direct answer and CoT accuracies for each reasoning category across models.
Figure 9: Performance of zero-shot direct (blue) and zero-shot CoT (orange) across datasets and models. Graphs are sorted in ascending order by median delta (CoT, direct). The datasets benefiting substantially are all symbolic or semi-symbolic in nature.
Table 7: Zero-shot accuracy for direct answering and CoT prompts on all datasets
Dataset Type Model zero-shot CoT accuracy zero-shot DA accuracy
MuSR Team Allocations Soft Reasoning Llama 2 7b 34.8 37.2
MuSR Team Allocations Soft Reasoning Mistral 7b 38.8 46.8
MuSR Team Allocations Soft Reasoning Llama 3.1 8b 44.0 48.0
MuSR Team Allocations Soft Reasoning Llama 3.1 70b 65.2 66.8
MuSR Team Allocations Soft Reasoning Gemma 2 9b 47.2 44.8
MuSR Team Allocations Soft Reasoning Phi-3 Small 8k 47.2 61.6
MuSR Team Allocations Soft Reasoning Qwen 2 7b 42.0 49.6
MuSR Team Allocations Soft Reasoning Qwen 2 72b 58.0 66.8
MuSR Team Allocations Soft Reasoning GPT-4o Mini 61.2 58.4
MuSR Team Allocations Soft Reasoning Gpt-4o 64.0 63.6
MuSR Team Allocations Soft Reasoning Claude-3 Haiku 56.8 59.2
MuSR Team Allocations Soft Reasoning Claude-3.5 Sonnet 80.4 63.2
MuSR Team Allocations Soft Reasoning Gemini 1.5 Flash 48.8 55.2
MuSR Team Allocations Soft Reasoning Gemini 1.5 Pro 58.4 62.4
SiQA Commonsense Llama 2 7b 53.4 55.9
SiQA Commonsense Mistral 7b 35.9 33.5
SiQA Commonsense Llama 3.1 8b 73.5 73.5
SiQA Commonsense Llama 3.1 70b 78.7 80.9
SiQA Commonsense Gemma 2 9b 74.9 76.3
SiQA Commonsense Phi-3 Small 8k 38.0 40.4
SiQA Commonsense Qwen 2 7b 37.3 39.3
SiQA Commonsense Qwen 2 72b 80.5 80.4
SiQA Commonsense GPT-4o Mini 79.0 80.0
SiQA Commonsense Gpt-4o 81.9 81.5
SiQA Commonsense Claude-3 Haiku 75.4 74.8
SiQA Commonsense Claude-3.5 Sonnet 79.7 81.0
SiQA Commonsense Gemini 1.5 Flash 74.5 79.1
SiQA Commonsense Gemini 1.5 Pro 73.9 78.2
MuSiQue Soft Reasoning Llama 2 7b 40.1 36.1
MuSiQue Soft Reasoning Mistral 7b 47.3 47.2
MuSiQue Soft Reasoning Llama 3.1 8b 62.6 64.7
MuSiQue Soft Reasoning Llama 3.1 70b 74.0 72.2
MuSiQue Soft Reasoning Gemma 2 9b 67.7 68.7
MuSiQue Soft Reasoning Phi-3 Small 8k 58.3 64.3
MuSiQue Soft Reasoning Qwen 2 7b 60.7 65.1
MuSiQue Soft Reasoning Qwen 2 72b 56.3 69.0
MuSiQue Soft Reasoning GPT-4o Mini 71.3 68.2
MuSiQue Soft Reasoning Gpt-4o 73.5 70.1
MuSiQue Soft Reasoning Claude-3 Haiku 54.8 56.0
MuSiQue Soft Reasoning Claude-3.5 Sonnet 66.9 70.4
MuSiQue Soft Reasoning Gemini 1.5 Flash 69.8 66.2
MuSiQue Soft Reasoning Gemini 1.5 Pro 69.8 71.3
AGIEval LSAT RC Soft Reasoning Llama 2 7b 31.2 36.4
AGIEval LSAT RC Soft Reasoning Mistral 7b 61.7 61.0
AGIEval LSAT RC Soft Reasoning Llama 3.1 8b 71.0 68.8
AGIEval LSAT RC Soft Reasoning Llama 3.1 70b 84.4 87.0
AGIEval LSAT RC Soft Reasoning Gemma 2 9b 75.1 78.1
AGIEval LSAT RC Soft Reasoning Phi-3 Small 8k 68.8 69.9
AGIEval LSAT RC Soft Reasoning Qwen 2 7b 61.0 66.5
AGIEval LSAT RC Soft Reasoning Qwen 2 72b 83.6 84.4
AGIEval LSAT RC Soft Reasoning GPT-4o Mini 77.3 74.3
AGIEval LSAT RC Soft Reasoning Gpt-4o 88.1 81.4
AGIEval LSAT RC Soft Reasoning Claude-3 Haiku 71.7 65.1
AGIEval LSAT RC Soft Reasoning Claude-3.5 Sonnet 90.0 89.6
AGIEval LSAT RC Soft Reasoning Gemini 1.5 Flash 78.1 81.0
AGIEval LSAT RC Soft Reasoning Gemini 1.5 Pro 82.2 85.9
CommonsenseQA Commonsense Llama 2 7b 49.4 54.6
CommonsenseQA Commonsense Mistral 7b 68.0 68.0
CommonsenseQA Commonsense Llama 3.1 8b 68.5 74.9
CommonsenseQA Commonsense Llama 3.1 70b 83.5 84.4
CommonsenseQA Commonsense Gemma 2 9b 79.2 80.1
CommonsenseQA Commonsense Phi-3 Small 8k 81.8 80.3
CommonsenseQA Commonsense Qwen 2 7b 78.5 79.0
CommonsenseQA Commonsense Qwen 2 72b 87.4 87.3
CommonsenseQA Commonsense GPT-4o Mini 82.5 83.9
CommonsenseQA Commonsense Gpt-4o 86.5 87.3
CommonsenseQA Commonsense Claude-3 Haiku 80.6 79.0
CommonsenseQA Commonsense Claude-3.5 Sonnet 85.1 84.3
CommonsenseQA Commonsense Gemini 1.5 Flash 79.7 82.6
CommonsenseQA Commonsense Gemini 1.5 Pro 79.9 82.9
GPQA Mathematical Llama 2 7b 28.3 24.3
GPQA Mathematical Mistral 7b 23.0 24.3
GPQA Mathematical Llama 3.1 8b 24.1 25.9
GPQA Mathematical Llama 3.1 70b 23.2 25.9
GPQA Mathematical Gemma 2 9b 26.3 21.2
GPQA Mathematical Phi-3 Small 8k 22.3 20.8
GPQA Mathematical Qwen 2 7b 24.1 24.6
GPQA Mathematical Qwen 2 72b 21.0 18.1
GPQA Mathematical GPT-4o Mini 21.0 24.0
GPQA Mathematical Gpt-4o 23.7 25.9
GPQA Mathematical Claude-3 Haiku 25.4 22.3
GPQA Mathematical Claude-3.5 Sonnet 25.4 25.9
GPQA Mathematical Gemini 1.5 Flash 22.3 22.8
GPQA Mathematical Gemini 1.5 Pro 21.0 23.7
AGIEval LSAT LR Soft Reasoning Llama 2 7b 29.4 33.5
AGIEval LSAT LR Soft Reasoning Mistral 7b 44.1 47.8
AGIEval LSAT LR Soft Reasoning Llama 3.1 8b 59.0 53.9
AGIEval LSAT LR Soft Reasoning Llama 3.1 70b 81.4 81.0
AGIEval LSAT LR Soft Reasoning Gemma 2 9b 64.9 67.6
AGIEval LSAT LR Soft Reasoning Phi-3 Small 8k 64.5 64.1
AGIEval LSAT LR Soft Reasoning Qwen 2 7b 50.6 58.4
AGIEval LSAT LR Soft Reasoning Qwen 2 72b 77.3 75.1
AGIEval LSAT LR Soft Reasoning GPT-4o Mini 65.3 68.2
AGIEval LSAT LR Soft Reasoning Gpt-4o 87.3 83.9
AGIEval LSAT LR Soft Reasoning Claude-3 Haiku 55.7 54.7
AGIEval LSAT LR Soft Reasoning Claude-3.5 Sonnet 83.7 82.7
AGIEval LSAT LR Soft Reasoning Gemini 1.5 Flash 70.0 71.2
AGIEval LSAT LR Soft Reasoning Gemini 1.5 Pro 79.4 80.4
PiQA Commonsense Llama 2 7b 62.1 64.7
PiQA Commonsense Mistral 7b 78.6 77.7
PiQA Commonsense Llama 3.1 8b 85.0 84.2
PiQA Commonsense Llama 3.1 70b 91.8 90.6
PiQA Commonsense Gemma 2 9b 84.0 84.8
PiQA Commonsense Phi-3 Small 8k 89.1 85.5
PiQA Commonsense Qwen 2 7b 84.3 86.2
PiQA Commonsense Qwen 2 72b 92.9 89.1
PiQA Commonsense GPT-4o Mini 93.1 88.6
PiQA Commonsense Gpt-4o 95.9 95.5
PiQA Commonsense Claude-3 Haiku 85.9 86.6
PiQA Commonsense Claude-3.5 Sonnet 94.6 94.5
PiQA Commonsense Gemini 1.5 Flash 84.6 89.8
PiQA Commonsense Gemini 1.5 Pro 88.1 91.3
Arc Easy Knowledge Llama 2 7b 71.1 69.8
Arc Easy Knowledge Mistral 7b 87.5 86.7
Arc Easy Knowledge Llama 3.1 8b 93.0 92.5
Arc Easy Knowledge Llama 3.1 70b 97.5 97.9
Arc Easy Knowledge Gemma 2 9b 94.9 95.8
Arc Easy Knowledge Phi-3 Small 8k 96.0 96.3
Arc Easy Knowledge Qwen 2 7b 89.5 84.7
Arc Easy Knowledge Qwen 2 72b 97.9 97.4
Arc Easy Knowledge GPT-4o Mini 96.8 94.6
Arc Easy Knowledge Gpt-4o 98.9 98.1
Arc Easy Knowledge Claude-3 Haiku 95.1 95.4
Arc Easy Knowledge Claude-3.5 Sonnet 98.6 98.4
Arc Easy Knowledge Gemini 1.5 Flash 96.8 97.2
Arc Easy Knowledge Gemini 1.5 Pro 97.2 94.6
Arc Challenge Knowledge Llama 2 7b 49.2 45.2
Arc Challenge Knowledge Mistral 7b 78.3 76.6
Arc Challenge Knowledge Llama 3.1 8b 86.0 82.6
Arc Challenge Knowledge Llama 3.1 70b 95.0 93.6
Arc Challenge Knowledge Gemma 2 9b 91.0 89.6
Arc Challenge Knowledge Phi-3 Small 8k 91.6 91.0
Arc Challenge Knowledge Qwen 2 7b 83.9 75.3
Arc Challenge Knowledge Qwen 2 72b 96.3 94.6
Arc Challenge Knowledge GPT-4o Mini 93.3 82.6
Arc Challenge Knowledge Gpt-4o 96.0 95.3
Arc Challenge Knowledge Claude-3 Haiku 89.3 89.3
Arc Challenge Knowledge Claude-3.5 Sonnet 96.0 95.3
Arc Challenge Knowledge Gemini 1.5 Flash 92.3 93.6
Arc Challenge Knowledge Gemini 1.5 Pro 91.6 90.6
AGIEval LSAT AR Soft Reasoning Llama 2 7b 17.0 17.4
AGIEval LSAT AR Soft Reasoning Mistral 7b 21.7 19.1
AGIEval LSAT AR Soft Reasoning Llama 3.1 8b 20.4 26.1
AGIEval LSAT AR Soft Reasoning Llama 3.1 70b 32.6 28.7
AGIEval LSAT AR Soft Reasoning Gemma 2 9b 24.8 23.0
AGIEval LSAT AR Soft Reasoning Phi-3 Small 8k 28.3 26.5
AGIEval LSAT AR Soft Reasoning Qwen 2 7b 27.0 23.9
AGIEval LSAT AR Soft Reasoning Qwen 2 72b 29.1 28.3
AGIEval LSAT AR Soft Reasoning GPT-4o Mini 32.2 23.0
AGIEval LSAT AR Soft Reasoning Gpt-4o 37.8 30.0
AGIEval LSAT AR Soft Reasoning Claude-3 Haiku 24.8 23.5
AGIEval LSAT AR Soft Reasoning Claude-3.5 Sonnet 38.3 33.9
AGIEval LSAT AR Soft Reasoning Gemini 1.5 Flash 27.8 27.8
AGIEval LSAT AR Soft Reasoning Gemini 1.5 Pro 30.0 31.7
BiGGen Bench Soft Reasoning Llama 2 7b 61.6 56.8
BiGGen Bench Soft Reasoning Mistral 7b 70.1 68.1
BiGGen Bench Soft Reasoning Llama 3.1 8b 66.5 67.7
BiGGen Bench Soft Reasoning Llama 3.1 70b 78.9 76.9
BiGGen Bench Soft Reasoning Gemma 2 9b 64.7 64.5
BiGGen Bench Soft Reasoning Phi-3 Small 8k 69.7 63.0
BiGGen Bench Soft Reasoning Qwen 2 7b 46.2 69.9
BiGGen Bench Soft Reasoning Qwen 2 72b 74.3 79.9
BiGGen Bench Soft Reasoning GPT-4o Mini 70.3 77.7
BiGGen Bench Soft Reasoning Gpt-4o 86.0 82.0
BiGGen Bench Soft Reasoning Claude-3 Haiku 80.0 80.0
BiGGen Bench Soft Reasoning Claude-3.5 Sonnet 91.4 79.3
BiGGen Bench Soft Reasoning Gemini 1.5 Flash 73.9 68.5
BiGGen Bench Soft Reasoning Gemini 1.5 Pro 78.7 67.1
Winogrande Commonsense Llama 2 7b 49.9 50.4
Winogrande Commonsense Mistral 7b 60.4 56.5
Winogrande Commonsense Llama 3.1 8b 66.5 63.3
Winogrande Commonsense Llama 3.1 70b 84.2 81.2
Winogrande Commonsense Gemma 2 9b 68.7 67.7
Winogrande Commonsense Phi-3 Small 8k 81.5 81.6
Winogrande Commonsense Qwen 2 7b 67.1 60.7
Winogrande Commonsense Qwen 2 72b 81.9 80.7
Winogrande Commonsense GPT-4o Mini 79.2 71.9
Winogrande Commonsense Gpt-4o 89.7 86.5
Winogrande Commonsense Claude-3 Haiku 70.7 66.2
Winogrande Commonsense Claude-3.5 Sonnet 89.4 85.7
Winogrande Commonsense Gemini 1.5 Flash 72.5 74.8
Winogrande Commonsense Gemini 1.5 Pro 75.5 78.3
MMLU Knowledge Llama 2 7b 46.3 41.7
MMLU Knowledge Mistral 7b 60.5 56.5
MMLU Knowledge Llama 3.1 8b 72.6 67.5
MMLU Knowledge Llama 3.1 70b 85.0 83.2
MMLU Knowledge Gemma 2 9b 73.8 71.4
MMLU Knowledge Phi-3 Small 8k 76.3 73.6
MMLU Knowledge Qwen 2 7b 67.0 64.5
MMLU Knowledge Qwen 2 72b 81.3 77.8
MMLU Knowledge GPT-4o Mini 79.9 74.8
MMLU Knowledge Gpt-4o 87.5 83.4
MMLU Knowledge Claude-3 Haiku 72.2 68.4
MMLU Knowledge Claude-3.5 Sonnet 87.2 84.0
MMLU Knowledge Gemini 1.5 Flash 76.3 74.7
MMLU Knowledge Gemini 1.5 Pro 81.3 81.1
StrategyQA Commonsense Llama 2 7b 39.5 31.2
StrategyQA Commonsense Mistral 7b 66.1 55.8
StrategyQA Commonsense Llama 3.1 8b 73.7 68.6
StrategyQA Commonsense Llama 3.1 70b 85.3 83.8
StrategyQA Commonsense Gemma 2 9b 73.7 66.4
StrategyQA Commonsense Phi-3 Small 8k 72.3 66.0
StrategyQA Commonsense Qwen 2 7b 63.2 54.8
StrategyQA Commonsense Qwen 2 72b 81.7 76.9
StrategyQA Commonsense GPT-4o Mini 82.2 84.5
StrategyQA Commonsense Gpt-4o 84.5 85.5
StrategyQA Commonsense Claude-3 Haiku 73.4 65.0
StrategyQA Commonsense Claude-3.5 Sonnet 80.1 76.3
StrategyQA Commonsense Gemini 1.5 Flash 72.5 75.2
StrategyQA Commonsense Gemini 1.5 Pro 74.0 71.4
MuSR Object Placements Soft Reasoning Llama 2 7b 36.3 30.5
MuSR Object Placements Soft Reasoning Mistral 7b 50.8 43.4
MuSR Object Placements Soft Reasoning Llama 3.1 8b 55.5 53.5
MuSR Object Placements Soft Reasoning Llama 3.1 70b 65.6 43.8
MuSR Object Placements Soft Reasoning Gemma 2 9b 63.3 57.0
MuSR Object Placements Soft Reasoning Phi-3 Small 8k 53.1 55.1
MuSR Object Placements Soft Reasoning Qwen 2 7b 48.8 48.4
MuSR Object Placements Soft Reasoning Qwen 2 72b 61.7 45.7
MuSR Object Placements Soft Reasoning GPT-4o Mini 59.0 55.0
MuSR Object Placements Soft Reasoning Gpt-4o 67.6 45.3
MuSR Object Placements Soft Reasoning Claude-3 Haiku 46.9 52.3
MuSR Object Placements Soft Reasoning Claude-3.5 Sonnet 69.5 51.2
MuSR Object Placements Soft Reasoning Gemini 1.5 Flash 61.7 56.2
MuSR Object Placements Soft Reasoning Gemini 1.5 Pro 66.4 50.0
FOLIO Symbolic Llama 2 7b 36.5 33.0
FOLIO Symbolic Mistral 7b 50.7 41.9
FOLIO Symbolic Llama 3.1 8b 58.6 56.7
FOLIO Symbolic Llama 3.1 70b 70.9 69.0
FOLIO Symbolic Gemma 2 9b 66.0 55.7
FOLIO Symbolic Phi-3 Small 8k 68.0 59.6
FOLIO Symbolic Qwen 2 7b 60.6 51.2
FOLIO Symbolic Qwen 2 72b 65.0 65.0
FOLIO Symbolic GPT-4o Mini 65.0 58.1
FOLIO Symbolic Gpt-4o 79.8 62.6
FOLIO Symbolic Claude-3 Haiku 61.6 48.8
FOLIO Symbolic Claude-3.5 Sonnet 73.9 68.5
FOLIO Symbolic Gemini 1.5 Flash 74.9 69.5
FOLIO Symbolic Gemini 1.5 Pro 73.9 74.4
ContextHub Deductive L2 Symbolic Llama 2 7b 34.8 12.6
ContextHub Deductive L2 Symbolic Mistral 7b 48.8 55.1
ContextHub Deductive L2 Symbolic Llama 3.1 8b 52.8 21.5
ContextHub Deductive L2 Symbolic Llama 3.1 70b 50.0 41.1
ContextHub Deductive L2 Symbolic Gemma 2 9b 50.0 43.0
ContextHub Deductive L2 Symbolic Phi-3 Small 8k 52.4 49.1
ContextHub Deductive L2 Symbolic Qwen 2 7b 51.3 39.8
ContextHub Deductive L2 Symbolic Qwen 2 72b 52.8 44.0
ContextHub Deductive L2 Symbolic GPT-4o Mini 47.0 42.0
ContextHub Deductive L2 Symbolic Gpt-4o 54.5 45.6
ContextHub Deductive L2 Symbolic Claude-3 Haiku 45.2 41.8
ContextHub Deductive L2 Symbolic Claude-3.5 Sonnet 53.0 46.2
ContextHub Deductive L2 Symbolic Gemini 1.5 Flash 45.0 39.5
ContextHub Deductive L2 Symbolic Gemini 1.5 Pro 57.3 43.3
ContextHub Abductive L2 Symbolic Llama 2 7b 34.3 31.9
ContextHub Abductive L2 Symbolic Mistral 7b 34.0 25.7
ContextHub Abductive L2 Symbolic Llama 3.1 8b 41.3 37.3
ContextHub Abductive L2 Symbolic Llama 3.1 70b 51.0 44.4
ContextHub Abductive L2 Symbolic Gemma 2 9b 41.5 32.9
ContextHub Abductive L2 Symbolic Phi-3 Small 8k 44.3 32.8
ContextHub Abductive L2 Symbolic Qwen 2 7b 37.8 33.4
ContextHub Abductive L2 Symbolic Qwen 2 72b 45.5 32.2
ContextHub Abductive L2 Symbolic GPT-4o Mini 65.0 55.0
ContextHub Abductive L2 Symbolic Gpt-4o 57.5 46.8
ContextHub Abductive L2 Symbolic Claude-3 Haiku 37.0 31.4
ContextHub Abductive L2 Symbolic Claude-3.5 Sonnet 56.8 40.4
ContextHub Abductive L2 Symbolic Gemini 1.5 Flash 53.1 32.2
ContextHub Abductive L2 Symbolic Gemini 1.5 Pro 53.5 43.7
MMLU Pro Knowledge Llama 2 7b 19.9 19.6
MMLU Pro Knowledge Mistral 7b 31.6 28.4
MMLU Pro Knowledge Llama 3.1 8b 44.8 38.0
MMLU Pro Knowledge Llama 3.1 70b 64.9 55.0
MMLU Pro Knowledge Gemma 2 9b 48.1 42.7
MMLU Pro Knowledge Phi-3 Small 8k 54.8 43.7
MMLU Pro Knowledge Qwen 2 7b 45.0 36.2
MMLU Pro Knowledge Qwen 2 72b 62.8 44.3
MMLU Pro Knowledge GPT-4o Mini 62.3 42.6
MMLU Pro Knowledge Gpt-4o 72.1 55.0
MMLU Pro Knowledge Claude-3 Haiku 47.6 39.0
MMLU Pro Knowledge Claude-3.5 Sonnet 73.4 57.2
MMLU Pro Knowledge Gemini 1.5 Flash 58.5 47.2
MMLU Pro Knowledge Gemini 1.5 Pro 65.3 57.4
MuSR Murder Mysteries Soft Reasoning Llama 2 7b 50.0 50.0
MuSR Murder Mysteries Soft Reasoning Mistral 7b 62.8 55.6
MuSR Murder Mysteries Soft Reasoning Llama 3.1 8b 70.4 57.2
MuSR Murder Mysteries Soft Reasoning Llama 3.1 70b 73.6 69.6
MuSR Murder Mysteries Soft Reasoning Gemma 2 9b 76.8 61.6
MuSR Murder Mysteries Soft Reasoning Phi-3 Small 8k 61.6 58.8
MuSR Murder Mysteries Soft Reasoning Qwen 2 7b 59.2 53.2
MuSR Murder Mysteries Soft Reasoning Qwen 2 72b 80.8 64.4
MuSR Murder Mysteries Soft Reasoning GPT-4o Mini 71.2 63.6
MuSR Murder Mysteries Soft Reasoning Gpt-4o 87.6 70.8
MuSR Murder Mysteries Soft Reasoning Claude-3 Haiku 62.4 56.8
MuSR Murder Mysteries Soft Reasoning Claude-3.5 Sonnet 85.2 70.4
MuSR Murder Mysteries Soft Reasoning Gemini 1.5 Flash 70.8 58.4
MuSR Murder Mysteries Soft Reasoning Gemini 1.5 Pro 77.6 64.0
ContextHub Deductive L1 Symbolic Llama 2 7b 47.7 08.3
ContextHub Deductive L1 Symbolic Mistral 7b 50.3 67.3
ContextHub Deductive L1 Symbolic Llama 3.1 8b 50.7 23.3
ContextHub Deductive L1 Symbolic Llama 3.1 70b 53.8 40.7
ContextHub Deductive L1 Symbolic Gemma 2 9b 56.3 39.2
ContextHub Deductive L1 Symbolic Phi-3 Small 8k 54.8 50.2
ContextHub Deductive L1 Symbolic Qwen 2 7b 59.3 43.3
ContextHub Deductive L1 Symbolic Qwen 2 72b 51.5 44.0
ContextHub Deductive L1 Symbolic GPT-4o Mini 49.3 41.5
ContextHub Deductive L1 Symbolic Gpt-4o 59.3 49.0
ContextHub Deductive L1 Symbolic Claude-3 Haiku 50.5 39.7
ContextHub Deductive L1 Symbolic Claude-3.5 Sonnet 54.5 47.0
ContextHub Deductive L1 Symbolic Gemini 1.5 Flash 47.3 38.5
ContextHub Deductive L1 Symbolic Gemini 1.5 Pro 57.3 46.0
ContextHub Abductive L1 Symbolic Llama 2 7b 29.4 16.4
ContextHub Abductive L1 Symbolic Mistral 7b 46.9 25.8
ContextHub Abductive L1 Symbolic Llama 3.1 8b 43.6 24.2
ContextHub Abductive L1 Symbolic Llama 3.1 70b 55.3 43.9
ContextHub Abductive L1 Symbolic Gemma 2 9b 61.9 58.9
ContextHub Abductive L1 Symbolic Phi-3 Small 8k 62.5 60.3
ContextHub Abductive L1 Symbolic Qwen 2 7b 52.2 47.5
ContextHub Abductive L1 Symbolic Qwen 2 72b 61.9 45.0
ContextHub Abductive L1 Symbolic GPT-4o Mini 61.1 42.2
ContextHub Abductive L1 Symbolic Gpt-4o 74.2 65.6
ContextHub Abductive L1 Symbolic Claude-3 Haiku 35.3 22.8
ContextHub Abductive L1 Symbolic Claude-3.5 Sonnet 80.8 60.3
ContextHub Abductive L1 Symbolic Gemini 1.5 Flash 66.4 47.2
ContextHub Abductive L1 Symbolic Gemini 1.5 Pro 62.2 60.0
Big-Bench Hard Symbolic Llama 2 7b 29.8 31.9
Big-Bench Hard Symbolic Mistral 7b 39.3 35.1
Big-Bench Hard Symbolic Llama 3.1 8b 62.8 45.6
Big-Bench Hard Symbolic Llama 3.1 70b 78.9 54.8
Big-Bench Hard Symbolic Gemma 2 9b 58.7 50.8
Big-Bench Hard Symbolic Phi-3 Small 8k 70.0 55.1
Big-Bench Hard Symbolic Qwen 2 7b 52.6 47.6
Big-Bench Hard Symbolic Qwen 2 72b 75.1 59.0
Big-Bench Hard Symbolic GPT-4o Mini 77.7 49.7
Big-Bench Hard Symbolic Gpt-4o 84.6 64.5
Big-Bench Hard Symbolic Claude-3 Haiku 62.4 47.3
Big-Bench Hard Symbolic Claude-3.5 Sonnet 83.6 56.9
Big-Bench Hard Symbolic Gemini 1.5 Flash 71.3 55.4
Big-Bench Hard Symbolic Gemini 1.5 Pro 71.6 50.3
MATH Mathematical Llama 2 7b 04.2 04.0
MATH Mathematical Mistral 7b 12.4 06.1
MATH Mathematical Llama 3.1 8b 47.2 13.8
MATH Mathematical Llama 3.1 70b 64.4 22.8
MATH Mathematical Gemma 2 9b 45.6 19.1
MATH Mathematical Phi-3 Small 8k 43.2 18.5
MATH Mathematical Qwen 2 7b 53.7 13.3
MATH Mathematical Qwen 2 72b 63.5 23.8
MATH Mathematical GPT-4o Mini 69.6 24.3
MATH Mathematical Gpt-4o 73.3 35.2
MATH Mathematical Claude-3 Haiku 32.7 17.4
MATH Mathematical Claude-3.5 Sonnet 63.8 34.6
MATH Mathematical Gemini 1.5 Flash 54.5 31.3
MATH Mathematical Gemini 1.5 Pro 62.1 39.4
GSM8k-Hard Mathematical Llama 2 7b 06.7 01.8
GSM8k-Hard Mathematical Mistral 7b 21.0 03.0
GSM8k-Hard Mathematical Llama 3.1 8b 34.4 06.0
GSM8k-Hard Mathematical Llama 3.1 70b 46.6 14.0
GSM8k-Hard Mathematical Gemma 2 9b 40.9 08.8
GSM8k-Hard Mathematical Phi-3 Small 8k 33.0 06.9
GSM8k-Hard Mathematical Qwen 2 7b 48.4 05.0
GSM8k-Hard Mathematical Qwen 2 72b 54.8 13.7
GSM8k-Hard Mathematical GPT-4o Mini 53.9 11.7
GSM8k-Hard Mathematical Gpt-4o 60.3 26.0
GSM8k-Hard Mathematical Claude-3 Haiku 45.3 09.6
GSM8k-Hard Mathematical Claude-3.5 Sonnet 50.8 32.3
GSM8k-Hard Mathematical Gemini 1.5 Flash 54.6 16.2
GSM8k-Hard Mathematical Gemini 1.5 Pro 58.2 26.2
GSM8k Mathematical Llama 2 7b 29.6 06.9
GSM8k Mathematical Mistral 7b 59.2 10.2
GSM8k Mathematical Llama 3.1 8b 85.4 18.5
GSM8k Mathematical Llama 3.1 70b 85.6 37.0
GSM8k Mathematical Gemma 2 9b 89.2 24.9
GSM8k Mathematical Phi-3 Small 8k 90.0 24.9
GSM8k Mathematical Qwen 2 7b 87.9 20.7
GSM8k Mathematical Qwen 2 72b 94.6 40.1
GSM8k Mathematical GPT-4o Mini 94.1 31.8
GSM8k Mathematical Gpt-4o 95.8 58.8
GSM8k Mathematical Claude-3 Haiku 89.4 22.9
GSM8k Mathematical Claude-3.5 Sonnet 96.1 62.2
GSM8k Mathematical Gemini 1.5 Flash 91.4 38.6
GSM8k Mathematical Gemini 1.5 Pro 92.7 52.4
F.2 Full Few-shot Results
Table 8: Few-shot accuracy for direct answering and CoT prompts on all datasets
Dataset Type Model few-shot CoT accuracy few-shot DA accuracy
AGIEval LSAT RC Soft Reasoning Llama 2 7b 33.1 38.7
AGIEval LSAT RC Soft Reasoning Mistral 7b 52.4 57.2
AGIEval LSAT RC Soft Reasoning Llama 3.1 8b 60.2 70.3
AGIEval LSAT RC Soft Reasoning Llama 3.1 70b 84.4 88.8
AGIEval LSAT RC Soft Reasoning Gemma 2 9b 74.3 79.2
AGIEval LSAT RC Soft Reasoning Phi-3 Small 8k 63.2 65.1
AGIEval LSAT RC Soft Reasoning Qwen 2 7b 61.7 68.8
AGIEval LSAT RC Soft Reasoning Qwen 2 72b 85.9 85.9
AGIEval LSAT RC Soft Reasoning GPT-4o Mini 77.3 71.4
AGIEval LSAT RC Soft Reasoning Gemini 1.5 Flash 79.2 81.8
AGIEval LSAT LR Soft Reasoning Llama 2 7b 33.7 34.7
AGIEval LSAT LR Soft Reasoning Mistral 7b 46.1 48.0
AGIEval LSAT LR Soft Reasoning Llama 3.1 8b 55.7 58.0
AGIEval LSAT LR Soft Reasoning Llama 3.1 70b 83.3 85.1
AGIEval LSAT LR Soft Reasoning Gemma 2 9b 65.7 68.2
AGIEval LSAT LR Soft Reasoning Phi-3 Small 8k 64.7 59.2
AGIEval LSAT LR Soft Reasoning Qwen 2 7b 54.1 61.2
AGIEval LSAT LR Soft Reasoning Qwen 2 72b 77.5 79.6
AGIEval LSAT LR Soft Reasoning GPT-4o Mini 68.4 64.5
AGIEval LSAT LR Soft Reasoning Gemini 1.5 Flash 68.6 72.9
GPQA Mathematical Mistral 7b 23.0 25.9
GPQA Mathematical Llama 3.1 8b 22.1 27.2
GPQA Mathematical Llama 3.1 70b 24.8 24.3
GPQA Mathematical Gemma 2 9b 19.9 22.3
GPQA Mathematical Phi-3 Small 8k 23.9 22.5
GPQA Mathematical Qwen 2 7b 23.4 21.2
GPQA Mathematical Qwen 2 72b 22.8 19.9
GPQA Mathematical GPT-4o Mini 20.0 20.0
GPQA Mathematical Gemini 1.5 Flash 21.9 24.6
CommonsenseQA Commonsense Llama 2 7b 18.2 19.2
CommonsenseQA Commonsense Mistral 7b 73.6 70.4
CommonsenseQA Commonsense Llama 3.1 8b 74.0 76.5
CommonsenseQA Commonsense Llama 3.1 70b 84.7 84.6
CommonsenseQA Commonsense Gemma 2 9b 81.8 80.8
CommonsenseQA Commonsense Phi-3 Small 8k 80.8 80.4
CommonsenseQA Commonsense Qwen 2 7b 80.3 72.9
CommonsenseQA Commonsense Qwen 2 72b 88.4 87.8
CommonsenseQA Commonsense GPT-4o Mini 84.7 84.7
CommonsenseQA Commonsense Gemini 1.5 Flash 81.7 83.3
AGIEval LSAT AR Soft Reasoning Llama 2 7b 19.6 18.7
AGIEval LSAT AR Soft Reasoning Mistral 7b 20.9 22.6
AGIEval LSAT AR Soft Reasoning Llama 3.1 8b 24.8 26.1
AGIEval LSAT AR Soft Reasoning Llama 3.1 70b 36.1 30.9
AGIEval LSAT AR Soft Reasoning Gemma 2 9b 22.2 28.7
AGIEval LSAT AR Soft Reasoning Phi-3 Small 8k 27.8 20.0
AGIEval LSAT AR Soft Reasoning Qwen 2 7b 24.3 23.0
AGIEval LSAT AR Soft Reasoning Qwen 2 72b 27.0 30.0
AGIEval LSAT AR Soft Reasoning GPT-4o Mini 28.7 26.1
AGIEval LSAT AR Soft Reasoning Gemini 1.5 Flash 28.3 20.4
MMLU Knowledge Llama 2 7b 49.0 42.8
MMLU Knowledge Mistral 7b 63.0 57.0
MMLU Knowledge Llama 3.1 8b 71.7 69.3
MMLU Knowledge Llama 3.1 70b 84.3 83.7
MMLU Knowledge Gemma 2 9b 74.7 72.4
MMLU Knowledge Phi-3 Small 8k 77.3 75.2
MMLU Knowledge Qwen 2 7b 69.9 68.6
MMLU Knowledge Qwen 2 72b 82.7 81.8
MMLU Knowledge GPT-4o Mini 82.3 77.8
MMLU Knowledge Gemini 1.5 Flash 78.1 79.0
StrategyQA Commonsense Llama 2 7b 57.9 30.9
StrategyQA Commonsense Mistral 7b 70.7 72.0
StrategyQA Commonsense Llama 3.1 8b 74.4 65.8
StrategyQA Commonsense Llama 3.1 70b 87.1 84.2
StrategyQA Commonsense Gemma 2 9b 77.1 73.3
StrategyQA Commonsense Phi-3 Small 8k 75.0 71.1
StrategyQA Commonsense Qwen 2 7b 71.9 58.9
StrategyQA Commonsense Qwen 2 72b 83.2 80.1
StrategyQA Commonsense GPT-4o Mini 83.0 86.2
StrategyQA Commonsense Gemini 1.5 Flash 77.0 80.3
ContextHub Abductive L2 Symbolic Llama 2 7b 36.2 35.0
ContextHub Abductive L2 Symbolic Mistral 7b 33.8 30.0
ContextHub Abductive L2 Symbolic Llama 3.1 8b 32.7 36.1
ContextHub Abductive L2 Symbolic Llama 3.1 70b 54.6 51.2
ContextHub Abductive L2 Symbolic Gemma 2 9b 44.8 33.2
ContextHub Abductive L2 Symbolic Phi-3 Small 8k 49.8 34.2
ContextHub Abductive L2 Symbolic Qwen 2 7b 39.6 35.0
ContextHub Abductive L2 Symbolic Qwen 2 72b 54.7 34.9
ContextHub Abductive L2 Symbolic GPT-4o Mini 62.0 60.0
ContextHub Abductive L2 Symbolic Gemini 1.5 Flash 48.6 47.8
ContextHub Abductive L1 Symbolic Llama 2 7b 21.4 16.7
ContextHub Abductive L1 Symbolic Mistral 7b 23.6 21.7
ContextHub Abductive L1 Symbolic Llama 3.1 8b 40.0 36.1
ContextHub Abductive L1 Symbolic Llama 3.1 70b 62.2 58.9
ContextHub Abductive L1 Symbolic Gemma 2 9b 48.9 59.4
ContextHub Abductive L1 Symbolic Phi-3 Small 8k 59.2 56.4
ContextHub Abductive L1 Symbolic Qwen 2 7b 48.6 38.9
ContextHub Abductive L1 Symbolic Qwen 2 72b 53.3 56.1
ContextHub Abductive L1 Symbolic GPT-4o Mini 77.2 59.2
ContextHub Abductive L1 Symbolic Gemini 1.5 Flash 79.7 68.6
MuSR Murder Mysteries Soft Reasoning Mistral 7b 62.0 56.4
MuSR Murder Mysteries Soft Reasoning Llama 3.1 8b 61.6 61.2
MuSR Murder Mysteries Soft Reasoning Llama 3.1 70b 73.2 68.0
MuSR Murder Mysteries Soft Reasoning Gemma 2 9b 81.6 62.0
MuSR Murder Mysteries Soft Reasoning Phi-3 Small 8k 62.0 53.6
MuSR Murder Mysteries Soft Reasoning Qwen 2 7b 56.0 55.6
MuSR Murder Mysteries Soft Reasoning Qwen 2 72b 80.4 66.0
MuSR Murder Mysteries Soft Reasoning GPT-4o Mini 76.0 69.6
MuSR Murder Mysteries Soft Reasoning Gemini 1.5 Flash 70.0 66.4
MuSR Team Allocations Soft Reasoning Mistral 7b 42.8 43.2
MuSR Team Allocations Soft Reasoning Llama 3.1 8b 59.6 51.6
MuSR Team Allocations Soft Reasoning Llama 3.1 70b 89.2 63.6
MuSR Team Allocations Soft Reasoning Gemma 2 9b 48.4 45.6
MuSR Team Allocations Soft Reasoning Phi-3 Small 8k 66.0 46.4
MuSR Team Allocations Soft Reasoning Qwen 2 7b 34.0 40.8
MuSR Team Allocations Soft Reasoning Qwen 2 72b 56.0 66.4
MuSR Team Allocations Soft Reasoning GPT-4o Mini 75.6 60.0
MuSR Team Allocations Soft Reasoning Gemini 1.5 Flash 90.0 54.4
MMLU Pro Knowledge Llama 2 7b 21.5 20.4
MMLU Pro Knowledge Mistral 7b 34.8 26.7
MMLU Pro Knowledge Llama 3.1 8b 44.7 38.0
MMLU Pro Knowledge Llama 3.1 70b 64.4 55.1
MMLU Pro Knowledge Gemma 2 9b 48.5 42.4
MMLU Pro Knowledge Phi-3 Small 8k 54.8 43.2
MMLU Pro Knowledge Qwen 2 7b 46.6 39.0
MMLU Pro Knowledge Qwen 2 72b 62.5 51.6
MMLU Pro Knowledge GPT-4o Mini 63.0 45.0
MMLU Pro Knowledge Gemini 1.5 Flash 59.4 50.6
MuSR Object Placements Soft Reasoning Mistral 7b 55.5 41.0
MuSR Object Placements Soft Reasoning Llama 3.1 8b 66.8 50.4
MuSR Object Placements Soft Reasoning Llama 3.1 70b 67.2 57.4
MuSR Object Placements Soft Reasoning Gemma 2 9b 68.0 58.2
MuSR Object Placements Soft Reasoning Phi-3 Small 8k 62.1 51.6
MuSR Object Placements Soft Reasoning Qwen 2 7b 46.9 43.8
MuSR Object Placements Soft Reasoning Qwen 2 72b 66.4 43.0
MuSR Object Placements Soft Reasoning GPT-4o Mini 67.0 47.0
MuSR Object Placements Soft Reasoning Gemini 1.5 Flash 73.0 54.7
ContextHub Deductive L2 Symbolic Llama 2 7b 34.7 15.0
ContextHub Deductive L2 Symbolic Mistral 7b 63.8 51.4
ContextHub Deductive L2 Symbolic Llama 3.1 8b 76.1 27.3
ContextHub Deductive L2 Symbolic Llama 3.1 70b 82.6 53.6
ContextHub Deductive L2 Symbolic Gemma 2 9b 61.9 47.6
ContextHub Deductive L2 Symbolic Phi-3 Small 8k 61.5 54.0
ContextHub Deductive L2 Symbolic Qwen 2 7b 55.3 36.4
ContextHub Deductive L2 Symbolic Qwen 2 72b 80.2 54.0
ContextHub Deductive L2 Symbolic GPT-4o Mini 59.0 41.0
ContextHub Deductive L2 Symbolic Gemini 1.5 Flash 90.2 42.5
ContextHub Deductive L1 Symbolic Llama 2 7b 34.7 16.0
ContextHub Deductive L1 Symbolic Mistral 7b 46.2 59.2
ContextHub Deductive L1 Symbolic Llama 3.1 8b 73.0 23.0
ContextHub Deductive L1 Symbolic Llama 3.1 70b 67.5 50.0
ContextHub Deductive L1 Symbolic Gemma 2 9b 66.0 45.7
ContextHub Deductive L1 Symbolic Phi-3 Small 8k 74.8 51.8
ContextHub Deductive L1 Symbolic Qwen 2 7b 58.8 37.5
ContextHub Deductive L1 Symbolic Qwen 2 72b 70.7 42.8
ContextHub Deductive L1 Symbolic GPT-4o Mini 59.2 44.3
ContextHub Deductive L1 Symbolic Gemini 1.5 Flash 89.3 49.8
MATH Mathematical Llama 2 7b 04.7 03.9
MATH Mathematical Mistral 7b 13.7 07.1
MATH Mathematical Llama 3.1 8b 41.2 14.2
MATH Mathematical Llama 3.1 70b 61.9 24.2
MATH Mathematical Gemma 2 9b 47.5 19.8
MATH Mathematical Phi-3 Small 8k 42.4 18.9
MATH Mathematical Qwen 2 7b 55.0 15.0
MATH Mathematical Qwen 2 72b 65.3 26.2
MATH Mathematical GPT-4o Mini 71.7 24.6
MATH Mathematical Gemini 1.5 Flash 54.7 32.3
GSM8K Mathematical Llama 2 7b 29.0 07.7
GSM8K Mathematical Mistral 7b 56.2 12.5
GSM8K Mathematical Llama 3.1 8b 86.4 20.1
GSM8K Mathematical Llama 3.1 70b 96.1 39.1
GSM8K Mathematical Gemma 2 9b 89.2 24.9
GSM8K Mathematical Phi-3 Small 8k 90.4 24.5
GSM8K Mathematical Qwen 2 7b 87.6 21.4
GSM8K Mathematical Qwen 2 72b 93.2 40.6
GSM8K Mathematical GPT-4o Mini 94.2 32.8
GSM8K Mathematical Gemini 1.5 Flash 90.6 40.4
F.3 Answer Extractor and Average Answer Span Results
In this section, we report the number of generations from each model on each dataset that our answer parser could not extract. “-1” denotes that a model was not run on a certain dataset due to context length limitations in the few-shot setting. We see that these unparseable rates are generally low across the board. The weakest models struggle on some of the most challenging datasets, but unparseable rates are all at or below 15%.
We also report the average character index of the beginning of the answer span that the answer parser extracted. Of particular note is that the direct answer prompts all return an answer within the first 60 characters, indicating that the answers are returned almost immediately, as desired. CoT completions are much longer.
Table 9: Percentage of responses per dataset per model that our answer parser could not extract an answer for in the zero-shot direct answer setting. Prompt modifications were made to decrease these numbers. No model is above 15%.
Zero-shot Direct Answer Unparseable Answer Rate by Percentage
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gpt-4o
Claude-3 Haiku
Claude-3.5 Sonnet
Gemini 1.5 Flash
Gemini 1.5 Pro
CommonsenseQA 01.9 02.5 01.1 00.0 00.8 00.1 01.6 00.7 00.0 00.0 00.1 00.0 00.1 00.2
StrategyQA 00.0 01.9 00.1 00.0 11.7 00.5 04.9 02.7 00.0 00.0 00.0 00.0 00.0 00.2
SiQA 00.2 06.6 00.0 00.1 03.9 00.3 00.1 03.0 00.1 00.1 00.0 00.0 00.0 00.4
PiQA 00.4 06.0 00.0 00.1 03.3 02.1 00.0 05.5 00.2 00.0 00.1 00.0 00.1 00.9
Winogrande 00.0 03.0 00.1 00.0 02.1 00.2 05.1 00.4 00.0 00.0 00.0 00.0 00.0 03.6
Arc Easy 00.0 01.8 00.5 00.0 00.0 00.2 09.1 00.7 03.5 00.4 00.2 00.0 00.0 03.2
Arc Challenge 00.0 02.3 01.0 00.0 00.3 00.7 10.7 00.7 10.0 00.7 00.0 00.0 00.0 05.0
AGIEval LSAT LR 00.4 00.0 00.0 00.0 00.0 00.2 00.0 00.0 00.0 02.5 00.0 00.0 00.2 00.2
AGIEval LSAT AR 00.4 00.0 00.0 00.0 04.3 03.9 00.0 00.0 00.0 08.7 00.0 00.0 00.0 00.0
AGIEval LSAT RC 00.4 00.4 00.0 00.0 00.0 00.0 00.0 00.0 00.0 09.7 00.0 00.0 00.4 00.4
ContextHub Deductive L1 00.0 00.0 00.0 00.0 01.2 00.0 02.3 00.0 00.0 00.0 00.2 00.0 00.0 00.2
ContextHub Deductive L2 00.0 00.0 00.0 00.0 00.0 00.0 02.2 01.0 00.0 00.0 02.8 00.0 00.0 00.0
ContextHub Abductive L1 00.0 00.0 00.0 00.0 00.0 00.3 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
ContextHub Abductive L2 00.0 00.0 00.0 00.0 00.0 00.1 01.5 00.2 00.0 00.0 00.8 00.0 00.0 00.0
MuSR Murder Mysteries 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
MuSR Team Allocations 00.0 00.0 00.0 00.0 03.6 00.0 00.0 00.0 00.0 00.0 00.0 00.0 08.4 00.4
MuSR Object Placements 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
MMLU 00.1 00.0 00.0 00.0 00.1 00.2 03.6 01.2 00.6 00.0 01.3 00.3 00.2 00.7
MMLU Pro 00.7 01.3 01.0 00.3 01.0 03.7 06.8 12.2 00.4 00.3 00.3 00.4 00.6 00.8
GPQA 01.3 07.1 00.0 00.0 08.7 12.7 05.4 15.2 00.0 00.0 01.6 00.0 00.0 00.7
MATH 00.6 06.9 00.3 00.2 00.1 00.1 03.5 03.0 00.8 00.0 00.3 00.0 00.4 00.6
GSM8k 00.2 04.1 02.5 00.0 02.7 00.0 01.7 00.2 00.0 00.0 12.7 05.5 00.0 00.0
BigGen Bench 04.6 00.3 00.9 00.1 00.5 01.0 01.3 01.0 01.3 00.0 00.0 00.1 00.4 00.3
GSM8k-Hard 04.8 07.6 02.0 02.4 00.4 00.2 03.2 01.1 00.1 00.5 05.2 00.5 00.2 00.0
MuSiQue 00.1 00.0 00.0 00.0 00.0 00.1 00.0 00.0 00.0 00.0 00.0 00.0 00.2 00.1
Folio 04.4 00.0 00.0 00.0 00.0 00.0 03.9 00.0 00.0 12.3 00.0 00.0 00.0 00.5
BigBench-Hard 00.0 00.0 00.0 07.4 00.0 00.2 00.0 00.0 00.0 00.3 00.0 04.5 00.2 12.8
Table 10: Percentage of responses per dataset per model that our answer parser could not extract an answer for in the zero-shot CoT setting. Prompt modifications were made to decrease these numbers. No model is above 15%.
Zero-shot CoT Unparseable Answer Rate by Percentage
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gpt-4o
Claude-3 Haiku
Claude-3.5 Sonnet
Gemini 1.5 Flash
Gemini 1.5 Pro
CommonsenseQA 02.9 01.3 08.6 00.0 00.6 00.1 00.0 00.0 01.6 00.0 00.2 00.3 02.4 02.6
StrategyQA 01.0 00.1 01.1 00.8 00.3 00.4 00.3 00.0 00.0 00.0 00.0 00.0 02.1 04.4
SiQA 00.8 01.8 00.3 00.1 01.6 00.0 00.1 00.1 00.0 00.0 00.3 00.1 03.5 04.0
PiQA 01.6 01.6 00.2 00.1 02.8 00.3 00.5 00.3 00.0 00.0 01.4 00.3 04.6 04.6
Winogrande 00.9 01.4 00.2 00.2 00.9 00.4 00.3 00.0 00.0 00.0 00.0 00.0 02.0 03.4
Arc Easy 00.2 00.4 00.2 00.0 00.5 01.6 01.6 00.0 00.5 00.0 00.0 00.0 00.4 00.5
Arc Challenge 00.0 00.7 00.0 00.0 00.0 00.0 00.7 00.0 00.0 00.0 00.0 00.0 00.7 00.7
AGIEval LSAT LR 03.3 02.2 00.0 00.0 01.2 00.0 02.0 00.0 00.0 00.0 00.0 00.0 00.8 00.2
AGIEval LSAT AR 04.8 07.0 06.1 02.2 05.7 05.2 04.3 00.4 01.3 01.3 00.0 00.4 04.8 01.7
AGIEval LSAT RC 07.1 01.1 00.0 00.0 00.7 03.0 06.7 00.0 00.0 00.0 00.0 00.0 00.0 00.4
ContextHub Deductive L1 00.7 01.3 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.3
ContextHub Deductive L2 00.2 00.4 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.5 00.4
ContextHub Abductive L1 00.6 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
ContextHub Abductive L2 00.0 00.2 00.1 00.0 00.0 00.3 00.0 00.0 00.0 00.0 00.0 00.0 00.5 00.4
MuSR Murder Mysteries 00.0 00.4 00.0 00.0 00.0 11.6 00.4 00.0 00.0 00.0 00.0 00.0 06.8 03.6
MuSR Team Allocations 05.2 03.2 00.8 00.0 00.8 00.4 00.4 00.0 00.0 00.0 00.0 00.0 00.0 00.0
MuSR Object Placements 00.0 01.6 00.0 00.0 00.4 00.8 00.0 00.0 00.0 00.0 00.0 00.0 02.0 00.4
MMLU 01.9 00.6 01.0 00.2 01.5 01.0 00.4 00.2 00.0 00.1 00.0 00.1 03.1 03.2
MMLU Pro 04.4 05.4 13.1 03.3 12.5 03.6 05.4 02.0 02.4 01.9 00.4 00.4 05.0 04.4
GPQA 04.5 10.3 09.4 01.6 08.5 01.8 03.8 00.7 00.0 00.0 00.0 00.0 11.8 15.0
MATH 01.6 05.5 08.2 02.5 02.3 01.6 03.0 00.4 00.4 00.5 00.9 00.0 01.7 01.0
GSM8k 01.7 01.4 00.7 10.5 00.4 00.6 00.4 00.0 00.0 00.0 00.3 00.0 00.1 00.1
BigGen Bench 05.0 00.4 00.5 00.1 00.5 00.4 00.3 09.5 00.0 00.0 00.0 00.1 00.4 00.1
GSM8k-Hard 02.1 08.7 10.2 04.5 10.7 03.2 03.5 01.0 00.8 00.5 03.0 01.8 00.4 02.7
MuSiQue 01.4 00.0 08.3 00.1 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.7 03.1
Folio 00.0 00.0 01.5 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 02.0 01.5
BigBench-Hard 03.8 05.4 01.8 00.4 01.3 00.1 00.4 00.3 00.0 00.0 00.0 00.0 01.2 00.9
Table 11: Percentage of responses per dataset per model that our answer parser could not extract an answer for in the few-shot direct answer setting. Prompt modifications were made to decrease these numbers. No model is above 15%.
Few-shot Direct Answer Unparseable Answer Rate by Percentage
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gemini 1.5 Flash
CommonsenseQA 00.0 00.1 00.2 00.0 01.3 00.9 09.9 01.3 00.0 00.6
AGIEval LSAT LR 06.7 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.2
AGIEval LSAT AR 02.6 00.0 00.0 00.0 03.5 05.2 00.0 00.0 00.0 00.0
AGIEval LSAT RC 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
ContextHub Deductive L1 00.0 02.8 00.0 00.0 00.0 10.7 00.3 00.0 00.0 00.0
ContextHub Deductive L2 00.0 00.1 00.0 00.0 00.0 00.3 00.2 00.0 00.0 00.0
ContextHub Abductive L1 00.0 02.8 00.0 00.0 00.0 00.6 00.0 00.0 00.0 00.0
ContextHub Abductive L2 00.0 02.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
MuSR Murder Mysteries -1.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.4
MuSR Team Allocations -1.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
MuSR Object Placements -1.0 00.0 00.0 00.0 00.4 01.2 00.0 00.0 00.0 00.0
MMLU 04.2 00.2 00.0 00.0 00.1 00.0 00.4 00.1 00.0 00.2
MMLU Pro 05.1 01.2 02.4 00.3 01.0 09.1 00.5 02.6 00.4 00.5
GPQA -1.0 01.3 00.0 00.0 03.6 07.4 13.4 01.1 00.0 00.0
MATH 00.3 05.9 00.3 00.2 00.1 00.1 01.6 02.2 00.0 00.3
GSM8k 00.1 00.1 00.5 00.0 00.1 02.2 00.0 00.2 00.0 00.0
Table 12: Percentage of responses per dataset per model that our answer parser could not extract an answer for in the few-shot CoT setting. Prompt modifications were made to decrease these numbers. No model is above 15%.
Few-shot CoT Unparseable Answer Rate by Percentage
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gemini 1.5 Flash
CommonsenseQA 00.7 00.9 01.8 00.1 00.2 00.1 00.0 00.0 00.0 03.4
AGIEval LSAT LR 00.6 00.8 00.4 00.0 01.4 03.1 00.8 00.0 00.0 00.6
AGIEval LSAT AR 02.2 09.1 03.9 00.9 11.7 03.0 03.5 01.7 00.0 01.3
AGIEval LSAT RC 07.8 05.9 00.0 00.0 01.9 09.3 02.6 00.0 00.0 02.2
ContextHub Deductive L1 00.2 00.0 00.2 00.0 00.0 00.0 00.0 00.0 00.0 00.3
ContextHub Deductive L2 00.9 00.0 00.2 00.0 00.0 00.0 00.0 00.0 00.0 00.3
ContextHub Abductive L1 00.8 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0 00.0
ContextHub Abductive L2 03.1 00.0 05.3 00.1 00.0 00.2 00.0 00.0 00.0 00.7
MuSR Murder Mysteries -1.0 01.2 00.0 00.0 00.4 00.8 00.0 00.0 00.0 14.0
MuSR Team Allocations -1.0 02.4 00.0 00.0 00.0 00.0 00.8 00.0 00.0 00.4
MuSR Object Placements -1.0 00.4 00.0 00.0 01.2 00.4 00.0 00.0 00.0 00.0
MMLU 00.6 00.8 01.1 00.2 01.5 00.7 00.3 00.2 00.2 02.5
MMLU Pro 00.6 01.9 08.5 02.1 14.1 01.8 01.9 00.8 01.1 03.9
GPQA -1.0 12.1 10.3 00.9 12.9 06.0 05.6 03.3 00.0 13.6
MATH 01.5 06.8 08.2 02.4 11.1 02.6 02.9 01.1 00.5 01.8
GSM8k 00.8 01.3 01.0 00.1 00.5 00.5 00.1 00.0 00.1 00.1
Table 13: Average character index of where the answer span begins in a generated response for each dataset and model pair for the zero-shot direct answer setting. We use these numbers as a proxy for the model following instructions (i.e. generating reasoning before an answer). Prompt modifications were made to ensure CoT prompts resulted in longer generations and direct answer prompts led to short generations.
Zero-shot Direct Answer Span Location By Character Index
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gpt-4o
Claude-3 Haiku
Claude-3.5 Sonnet
Gemini 1.5 Flash
Gemini 1.5 Pro
CommonsenseQA 09 08 27 08 08 08 10 08 08 10 07 07 08 08
StrategyQA 44 45 27 44 44 44 46 44 08 08 42 41 08 07
SiQA 08 08 08 08 08 08 29 08 08 08 06 06 08 08
PiQA 07 08 08 08 08 08 25 08 08 08 04 05 08 08
Winogrande 08 09 08 08 08 08 09 08 08 09 05 04 08 08
Arc Easy 09 08 08 08 08 08 09 08 08 08 07 07 08 08
Arc Challenge 08 08 08 08 08 08 09 08 08 08 07 07 08 08
AGIEval LSAT LR 25 24 24 24 24 24 25 24 43 21 25 25 26 26
AGIEval LSAT AR 25 24 24 24 24 24 26 24 48 23 25 25 26 27
AGIEval LSAT RC 25 24 24 24 24 24 25 24 31 18 25 25 26 25
ContextHub Deductive L1 19 19 19 19 20 19 19 19 19 19 20 20 19 20
ContextHub Deductive L2 19 19 19 19 19 19 19 19 19 19 20 20 19 19
ContextHub Abductive L1 19 19 19 19 20 19 19 19 19 19 20 20 19 19
ContextHub Abductive L2 19 19 19 19 20 19 19 19 19 19 20 20 19 19
MuSR Murder Mysteries 08 08 27 08 08 08 08 08 08 08 06 04 08 08
MuSR Team Allocations 27 22 19 19 27 23 26 22 08 08 30 20 08 08
MuSR Object Placements 08 08 27 08 08 08 08 08 08 08 07 06 08 08
MMLU 19 18 19 19 20 18 18 18 19 19 19 19 19 20
MMLU Pro 20 19 38 19 21 19 19 20 19 19 20 20 19 19
GPQA 19 19 19 19 21 19 19 19 19 19 20 20 19 20
MATH 30 31 28 28 28 30 30 33 28 28 31 29 28 28
GSM8k 22 29 30 28 28 37 24 28 28 28 29 28 28 28
GSM8k-Hard 09 57 11 21 09 13 40 20 07 08 08 08 08 08
Folio 39 08 08 08 08 08 31 13 08 16 05 06 08 70
BigBench-Hard 39 22 25 21 26 32 29 26 28 19 28 28 10 16
Table 14: Average character index of where the answer span begins in a generated response for each dataset and model pair for the zero-shot CoT setting. We use these numbers as a proxy for the model following instructions (i.e. generating reasoning before an answer). Prompt modifications were made to ensure CoT prompts resulted in longer generations and direct answer prompts led to short generations.
Zero-shot CoT Answer Span Location By Character Index
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gpt-4o
Claude-3 Haiku
Claude-3.5 Sonnet
Gemini 1.5 Flash
Gemini 1.5 Pro
CommonsenseQA 441 564 845 1237 236 466 577 341 899 1086 626 1103 214 165
StrategyQA 726 434 996 1131 267 460 363 358 692 1033 754 1158 256 195
SiQA 569 423 841 965 235 528 472 420 847 1094 602 1016 196 169
PiQA 699 455 869 914 207 532 447 364 683 935 578 1092 200 150
Winogrande 377 324 645 694 187 326 391 298 634 750 408 889 200 173
Arc Easy 684 581 1154 1319 367 610 534 355 990 1239 789 1222 340 231
Arc Challenge 763 644 1178 1316 422 596 571 387 1020 1269 828 1240 372 267
AGIEval LSAT LR 2053 1324 1163 1675 524 689 1560 768 949 998 1561 728 906 886
AGIEval LSAT AR 1377 1791 1422 2182 712 1027 1819 1264 1230 1151 1202 849 817 871
AGIEval LSAT RC 1977 1032 1103 1575 779 590 1170 660 973 1079 1628 786 703 709
ContextHub Deductive L1 694 368 759 711 383 327 539 402 540 580 542 556 320 254
ContextHub Deductive L2 842 472 1095 990 614 442 789 585 840 758 777 655 515 503
ContextHub Abductive L1 577 461 747 879 464 440 754 638 788 879 683 594 368 325
ContextHub Abductive L2 861 600 1270 1229 686 571 976 856 1115 1113 894 894 601 551
MuSR Murder Mysteries 495 1592 1958 1847 1210 1246 1241 1718 1961 1965 1671 1759 1349 1213
MuSR Team Allocations 1212 1845 2294 2310 1513 1433 2021 2213 2562 2698 1479 1856 1596 1607
MuSR Object Placements 917 625 1354 1266 695 641 904 819 1593 1536 1210 1455 616 429
MMLU 834 512 663 622 503 277 497 407 400 461 447 409 630 413
MMLU Pro 1371 513 788 716 640 518 954 699 926 940 590 653 660 774
GPQA 1034 778 917 901 806 500 1018 628 541 666 486 472 981 735
MATH 742 1118 1222 1179 748 670 1189 1145 1125 1153 677 675 679 698
GSM8k 572 637 834 719 453 521 709 645 1048 1035 708 680 541 437
GSM8k-Hard 916 939 1027 1069 555 766 1083 1053 1350 1266 594 815 605 512
Folio 724 765 1479 1379 733 668 919 488 1285 1583 907 1194 934 492
BigBench-Hard 596 230 876 861 429 349 315 443 877 973 545 863 455 346
Table 15: Average character index of where the answer span begins in a generated response for each dataset and model pair for the few-shot direct answer setting. We use these numbers as a proxy for the model following instructions (i.e. generating reasoning before an answer). Prompt modifications were made to ensure CoT prompts resulted in longer generations and direct answer prompts led to short generations.
Few-shot Direct Answer Span Location By Character Index
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gemini 1.5 Flash
CommonsenseQA 87 08 27 08 08 08 10 08 08 08
AGIEval LSAT LR 25 24 24 24 24 24 24 24 31 24
AGIEval LSAT AR 25 24 24 24 24 24 24 24 27 24
AGIEval LSAT RC 25 24 24 24 24 24 24 24 25 24
ContextHub Deductive L1 19 19 19 19 19 19 19 19 19 19
ContextHub Deductive L2 19 19 19 19 19 19 19 19 19 19
ContextHub Abductive L1 19 19 19 19 19 19 19 19 19 19
ContextHub Abductive L2 19 19 19 19 19 19 19 19 19 19
MuSR Murder Mysteries -1 08 27 08 08 08 08 08 08 08
MuSR Team Allocations -1 21 19 19 27 21 27 23 08 08
MuSR Object Placements -1 08 27 08 08 08 08 08 08 08
MMLU 19 18 19 19 19 18 18 18 19 19
MMLU Pro 19 19 38 19 20 20 19 19 19 19
GPQA -1 19 19 19 19 19 19 19 19 19
MATH 29 36 29 29 28 30 30 41 28 28
GSM8k 22 23 23 22 22 23 22 24 27 28
Table 16: Average character index of where the answer span begins in a generated response for each dataset and model pair for the few-shot CoT setting. We use these numbers as a proxy for the model following instructions (i.e. generating reasoning before an answer). Prompt modifications were made to ensure CoT prompts resulted in longer generations and direct answer prompts led to short generations.
Few-shot CoT Answer Span Location By Character Index
dataset
Meta-Llama 2 7b
Mistral 7b
Meta-Llama 3.1 8b
Meta-Llama 3.1 70b
Gemma 2 9b
Phi-3 Small 8k
Qwen 2 7b
Qwen 2 72b
GPT-4o Mini
Gemini 1.5 Flash
CommonsenseQA 301 195 470 921 145 192 280 174 219 158
AGIEval LSAT LR 1037 510 464 539 437 359 530 599 894 523
AGIEval LSAT AR 1024 1247 886 768 573 1025 750 835 1033 670
AGIEval LSAT RC 799 378 131 206 164 111 241 205 1086 266
ContextHub Deductive L1 383 386 406 376 359 376 388 364 416 366
ContextHub Deductive L2 736 767 829 822 823 855 612 807 884 809
ContextHub Abductive L1 301 386 428 450 431 413 541 447 575 379
ContextHub Abductive L2 709 586 967 754 804 784 829 821 905 815
MuSR Murder Mysteries -1 1280 1693 1702 1225 1338 1246 1719 1974 1419
MuSR Team Allocations -1 2195 2087 2160 1628 1755 2181 2156 2632 1841
MuSR Object Placements -1 907 1104 1213 706 919 676 963 1351 853
MMLU 282 266 333 245 265 260 267 243 392 218
MMLU Pro 429 397 424 411 516 425 541 325 681 396
GPQA -1 848 782 774 615 711 662 703 670 594
MATH 630 705 584 640 747 529 1074 848 1261 553
GSM8k 374 332 352 352 398 372 415 341 651 314
Appendix G Zoom-in: MMLU and MMLU Pro
Table 17: The top 3 slices benefiting the most from CoT across MMLU and MMLU Pro for Llama 3.1 8b and 70b. 6 out of 12 of these top slices directly contain “math” or “mathematics.” We dive deeper into each category subsequently and observe that the questions leading to improvements in the other categories are mathematical in nature as well.
MMLU MMLU Pro
Model Subject Direct (%) CoT (%) Err. Red. (%) N Subject Direct (%) CoT (%) Err. Red. (%) N
Llama 3.1 8b elementary_mathematics 46.8 88.4 78.1 378 math 23.6 44.8 27.8 1350
Llama 3.1 8b high_school_mathematics 39.6 71.5 52.8 270 business 29.4 45.6 23.0 789
Llama 3.1 8b miscellaneous 83.9 89.9 37.3 783 physics 27.9 41.4 18.8 1299
Llama 3.1 70b elementary_mathematics 82.3 94.7 70.1 378 math 44.5 68.3 42.9 1351
Llama 3.1 70b medical_genetics 93.0 97.0 57.1 100 business 44.0 67.8 42.5 789
Llama 3.1 70b high_school_mathematics 61.5 82.2 53.8 270 chemistry 40.5 64.0 39.6 1132
MMLU and MMLU Pro show gains from adding CoT, but because these datasets are so broad, they defy simple characterization. We explore the performance of CoT on each category of MMLU to understand divergences in CoT performance between these domains. We list the top three categories where CoT gives the largest error reduction for Llama 3.1 8B and 70B on MMLU and MMLU Pro in Table 17. Some of these categories are explicitly mathematical in nature, as we might expect from Figure 3. We can also see that CoT is helping on categories like “business”; upon closer inspection, we found that these categories frequently involve math as well (e.g., business questions may involve computations surrounding wealth). We need to more carefully characterize MMLU at the instance level. In doing so, we can test our hypotheses with much finer granularity than possible by relying on subjective groupings into tasks and categories.
Breakdown by the presence of equations
We aim to design an instance-level classifier to determine if CoT is expected to help on a question or not. That is, we want a function $g:\mathbf{q}\rightarrow\{0,1\}$ where $g(\mathbf{q})$ returns 1 if $\mathrm{extract}(\tilde{\mathbf{y}}_{cot})=\mathbf{y}^{*}$ and $\mathrm{extract}(\tilde{\mathbf{y}}_{da})\neq\mathbf{y}^{*}$ where $\mathbf{y}^{*}$ is the gold answer to $\mathbf{q}$. We explored different forms of $g$; however, we ultimately found it most effective to use a classifier $g:(\mathbf{q},\tilde{\mathbf{y}}_{\mathrm{cot}})\rightarrow\{0,1\}$ which also consults the chain-of-thought produced by the model. This allows us to featurize how the LM solves the problem, particularly whether it uses symbolic reasoning or not.
We find that $g$ can be implemented with a single feature: does $\mathbf{q}$ or $\tilde{\mathbf{y}}_{\mathrm{cot}}$ contain a “=”? The “=” token very strongly indicates the presence of equations in the problem or its solution, which turn out to be a strong hallmark of symbolic reasoning. [^6] We explored implementing $g$ with a logistic regression classifier with tf-idf features over the $(\mathbf{q},\tilde{\mathbf{y}}_{\mathrm{cot}})$ pairs, trained over a subset of the data from MMLU and MMLU Pro. This classifier actually allowed us to discover the “=” feature, but its accuracy did not exceed the accuracy of that single feature.
We plot the overall CoT delta (performance of CoT minus the performance of direct answer) for both MMLU and MMLU Pro across multiple models between two bins according to this classifier $g$, labeled as “With =” and “Without =”, in Figure 4. We also report the amount of performance gain explained by questions having an “=” vs. not in Appendix G.1. We find that the majority of the performance gain from CoT on MMLU and MMLU Pro comes from questions that have an “=” in the question or generated responses. Because “=” are usually found in math problems, we equate this to CoT primarily benefiting MMLU and MMLU Pro on the math-related questions with very little to no gain (depending on the model) for non-math questions.
G.1 Performance impacts of “=” on MMLU and MMLU Pro
Tables 18 and 19 show the amount of total improvement from using CoT over direct prompting that can be explained by the presence of “=” on MMLU and MMLU Pro over multiple models.
Model Total CoT Delta CoT delta w/ = CoT delta w/o = Perf. Gain w/ = Fraction of N w/ =
Llama 2 7b 06.0 00.6 05.4 09.8% 10.9%
Mistral 7b 04.1 01.2 02.9 28.6% 09.8%
Llama 3.1 8b 05.5 02.9 02.6 52.9% 09.6%
Llama 3.1 70b 01.9 01.8 00.1 94.0% 10.6%
Gemma 2 9b 02.6 02.0 00.6 78.5% 10.0%
Phi-3 Small 8k 03.1 01.5 01.7 47.4% 08.3%
Qwen 2 7b 02.5 03.0 -0.5 100.0%0 09.8%
Qwen 2 72b 03.5 02.4 01.1 67.8% 09.6%
GPT-4o Mini 05.2 03.5 01.7 66.9% 10.5%
GPT-4o 04.2 02.4 01.8 57.6% 10.3%
Claude-3 Haiku 03.7 02.4 01.3 64.4% 09.3%
Claude-3.5 Sonnet 03.2 02.3 00.9 72.1% 10.7%
Gemini 1.5 Flash 03.0 01.7 01.2 59.0% 10.1%
Gemini 1.5 Pro 01.9 01.0 00.9 51.9% 09.6%
Table 18: Total CoT deltas on MMLU broken down by the total gain from questions and responses with an “=” vs. without an “=”.
Model Total CoT Delta CoT delta w/ = CoT delta w/o = Perf. Gain w/ = Fraction of N w/ =
Llama 2 7b 01.6 01.3 00.3 79.6% 43.6%
Mistral 7b 03.8 01.9 01.9 50.7% 41.8%
Llama 3.1 8b 12.4 10.0 02.4 80.8% 35.2%
Llama 3.1 70b 11.4 11.1 00.3 97.6% 39.6%
Gemma 2 9b 07.6 07.4 00.2 97.9% 40.2%
Phi-3 Small 8k 11.6 09.9 01.7 85.7% 42.7%
Qwen 2 7b 10.0 08.9 01.1 88.6% 41.6%
Qwen 2 72b 19.0 16.1 02.9 84.7% 41.4%
GPT-4o Mini 20.6 18.4 02.3 89.0% 44.0%
GPT-4o 17.7 17.1 00.6 96.7% 44.1%
Claude-3 Haiku 08.7 07.8 00.9 90.1% 42.0%
Claude-3.5 Sonnet 16.2 14.8 01.3 91.9% 43.4%
Gemini 1.5 Flash 12.9 11.8 01.1 91.3% 42.3%
Gemini 1.5 Pro 10.0 08.6 01.4 85.7% 41.8%
Table 19: Total CoT deltas on MMLU Pro broken down by the total gain from questions and responses with an “=” vs. without an “=”.
Appendix H Full results of evaluations on formal reasoning datasets
Dataset Method Mistral 7b Llama 3.1 8b Llama 3.1 70b GPT-4o Mini
Acc. % Unp. Acc. % Unp. Acc. % Unp. Acc. % Unp.
GSM8K Direct Answer 12.5 0.1 20.1 0.5 39.1 0.0 32.8 0.0
GSM8K CoT 56.2 1.4 86.4 1.0 96.1 0.1 94.2 0.1
GSM8K Plan + CoT Solver 45.0 1.0 78.7 0.4 94.7 0.0 92.0 0.1
GSM8K Plan + Direct Solver 10.6 0.1 19.6 0.1 42.2 0.0 39.3 0.0
GSM8K Plan + Tool Solver 59.8 8.6 80.3 1.3 94.4 0.4 90.5 1.5
GSM8K-Hard Direct Answer 2.9 0.7 4.4 0.6 12.8 0.7 12.3 7.6
GSM8K-Hard CoT 20.3 5.0 32.4 9.6 47.8 4.4 52.2 0.5
GSM8K-Hard Plan + CoT Solver 18.7 2.6 32.4 1.3 49.7 0.6 51.5 0.3
GSM8K-Hard Plan + Direct Solver 3.0 0.5 5.5 0.8 15.8 0.1 17.4 0.3
GSM8K-Hard Plan + Tool Solver 44.2 8.9 57.9 1.2 68.0 0.5 70.4 1.4
ContextHub Deductive L1 Direct Answer 59.2 2.8 23.0 0.0 50.0 0.0 44.3 0.0
ContextHub Deductive L1 CoT 46.2 0.2 73.0 0.2 67.5 0.0 59.2 0.0
ContextHub Deductive L1 Plan + CoT Solver 49.5 0.0 64.8 0.0 65.5 0.0 63.2 0.0
ContextHub Deductive L1 Plan + Direct Solver 45.8 3.0 55.8 0.0 53.5 0.0 56.2 0.0
ContextHub Deductive L1 Plan + Tool Solver 68.8 27.8 84.2 11.8 91.7 9.8 90.7 7.8
ContextHub Abductive L1 Direct Answer 21.7 2.8 36.1 0.0 58.9 0.0 59.2 0.0
ContextHub Abductive L1 CoT 23.9 0.0 40.0 0.0 62.2 0.0 76.9 0.0
ContextHub Abductive L1 Plan + CoT Solver 38.3 0.0 42.5 0.0 65.6 0.0 74.2 0.0
ContextHub Abductive L1 Plan + Direct Solver 46.9 3.9 33.3 0.3 63.1 0.0 61.7 0.0
ContextHub Abductive L1 Plan + Tool Solver 59.2 35.8 70.8 9.7 73.9 4.2 74.7 10.3
FOLIO Direct Answer 56.2 12.3 59.6 0.0 69.5 0.0 64.0 0.0
FOLIO CoT 53.7 1.5 56.7 2.5 72.4 2.0 70.4 0.0
FOLIO Plan + CoT Solver 53.7 0.0 55.7 0.0 73.9 0.5 70.4 0.0
FOLIO Plan + Direct Solver 52.7 0.0 54.2 0.0 72.9 0.0 63.5 0.0
FOLIO Plan + Tool Solver 48.8 46.8 54.2 28.6 70.0 16.7 62.6 25.1
Table 20: Performance and unparseable rates for few-shot direct answer, few-shot CoT, Plan + Direct Solver, Plan + CoT Solver, and Plan + Tool Solver Solver. “Acc.” stands for accuracy and “Unp.” stands for the rate of unparseable model responses that either fail to pass our answer extraction parser (for all methods except Plan + Tool Solver prompting) or fail to be executed by symbolic solvers. For FOLIO and ContextHub, we compute the accuracy by making a random guess for the unparseable responses; for GSM8K and GSM8K-Hard, we consider the unparseable responses as incorrect.
As discussed in Section 5, we include detailed evaluation results of few-shot direct answer, few-shot CoT, direct answer solver, CoT solver, and tool-augmented prompting in Table 20. The unparseable rate stands for the rate of unparseable model responses that either fail to pass our answer extraction parser (for all methods except tool-augmented prompting) or fail to be executed by symbolic solvers. For FOLIO and ContextHub, we compute the accuracy by making a random guess for the unparseable responses; for GSM8K and GSM8K-Hard, we consider the unparseable responses as incorrect.
We note that all models have a low unparseable rate ($<10\%$) for all methods except tool-augmented prompting. By manually inspecting the outputs, we observe that the high unparseable rate for some models with tool-augmented prompting is caused by these models generating Python programs or formal specifications that fail to follow the format of the formal language (Python or z3) and that lead to execution errors. Such an issue is particularly severe for the smaller models. However, we note that despite the high unparseable rate, the overall accuracy of these models with tool augmentation is still on par with or outperforms other methods.
Appendix I Discussion of Limitations
I.1 Long Horizon Planning
One set of tasks where symbolic reasoning helps substantially that our experiments haven’t covered as thoroughly (with the exception of BiGGen-Bench) is long-horizon planning (Valmeekam et al., 2023; Xie et al., 2024; Gundawar et al., 2024; Valmeekam et al., 2024). There are two reasons we don’t treat it here. First, we are primarily interested in tasks that are conveyed in language, and we see less complex planning in language-only tasks. Second, there has already been a large debate on the effectiveness of CoT, both pro (Huang et al., 2022; Hu et al., 2023) and against (Valmeekam et al., 2023; Kambhampati, 2024; Kambhampati et al., 2024b; Stechly et al., 2024a; Guan et al., 2024; Verma et al., 2024; Gundawar et al., 2024; Stechly et al., 2024b) using CoT and its derivatives like tree-of-thought (Yao et al., 2023; Kang et al., 2024), that has resulted in complex systems to help solve planning problems better. While story generation and interpretation involve elements of planning with natural language (Peng et al., 2022; Karpinska et al., 2024), such tasks are not conventionally formalized and benchmarked as planning and reasoning.
I.2 Dataset contamination
One limitation of our study is the presence of possible data contamination: it is unknown which benchmarks may have been explicitly pre-trained on by language models. If a model had memorized answers to benchmark questions, we would expect direct answering to close some of the gap with CoT, as the model can just reproduce a known answer rather than deriving it from scratch. We argue there are four reasons that our general conclusions are still trustworthy. First, we use a range of language model scales, including small models that have less capacity to memorize. Second, datasets with poor direct answering performance like GSM8K-Hard are unlikely to have been substantially memorized. Third, the inclusion of recent datasets such as MuSR (Sprague et al., 2024) and BiGGen Bench (Kim et al., 2024) helps to defray this risk. Fourth, our survey of the literature includes papers that were submitted to conferences in 2023, representing a range of older LLMs trained at various times.
Appendix J Example prompts
We will release all prompts and model responses on our Huggingface repo. We list a few prompt response pairs here in this section.
[…]
AGIEval LSAT AR zero-shot CoT response for Llama 3.1 70B To determine which one of the following could be the schedule of the students’ reports, let’s analyze each answer choice in light of the given conditions. First, let’s recall the conditions: 1. Exactly six students will give reports over three days. 2. Tuesday is the only day George can give a report. 3. Neither Olivia nor Robert can give an afternoon report. 4. If Nina gives a report, then on the next day Helen and Irving must both give reports, unless Nina’s report is given on Wednesday. Let’s evaluate each answer choice: (A) Mon. morning: Helen; Mon. afternoon: Robert Tues. morning: Olivia; Tues. afternoon: Irving Wed. morning: Lenore; Wed. afternoon: Kyle - This choice does not violate any conditions directly. However, we need to check if it aligns with all the rules, especially considering the Nina rule, which isn’t directly applicable here since Nina isn’t scheduled. George’s requirement is also not met since George isn’t scheduled on Tuesday. (B) Mon. morning: Irving; Mon. afternoon: Olivia Tues. morning: Helen; Tues. afternoon: Kyle Wed. morning: Nina; Wed. afternoon: Lenore - This choice violates the condition that if Nina gives a report, then on the next day Helen and Irving must both give reports, unless Nina’s report is given on Wednesday. Since Nina’s report is on Wednesday, this condition is technically not violated, but we need to ensure all other conditions are met. George is not scheduled on Tuesday, which is a requirement. (C) Mon. morning: Lenore; Mon. afternoon: Helen Tues. morning: George; Tues. afternoon: Kyle Wed. morning: Robert; Wed. afternoon: Irving - This choice meets the condition that George gives a report on Tuesday. It also does not violate any other conditions directly. However, we need to ensure that the Nina rule isn’t applicable or is met, which in this case, it’s not applicable since Nina isn’t scheduled. (D) Mon. morning: Nina; Mon. afternoon: Helen Tues. morning: Robert; Tues. afternoon: Irving Wed. morning: Olivia; Wed. afternoon: Lenore - This choice violates the condition that if Nina gives a report, then on the next day Helen and Irving must both give reports. Here, Helen is already scheduled on the same day as Nina, and while Irving is scheduled the next day, the condition specifies both Helen and Irving must give reports the next day, which is not the case since Helen’s report is on the same day as Nina’s. (E) Mon. morning: Olivia; Mon. afternoon: Nina Tues. morning: Irving; Tues. afternoon: Helen Wed. morning: Kyle; Wed. afternoon: George - This choice violates the condition that Tuesday is the only day George can give a report, as George is scheduled on Wednesday. Given the analysis, the only choice that does not directly violate any of the given conditions and meets the specific requirements (especially considering the placement of George and the handling of the Nina rule where applicable) is: The answer is therefore (C).
[…]
CommonsenseQA zero-shot direct answer prompt for Llama 3.1 70B <|start_header_id|>system<|end_header_id|> You are a helpful AI assistant that will answer reasoning questions. You will always say at the end \"Answer: <Your Answer Letter Choice>\". You must only pick one answer and you must end your response with \"Answer: <Your Answer Letter Choice>\" everytime!<|eot_id|><|start_header_id|>user<|end_header_id|> Question: Where would you find magazines along side many other printed works? Answer Choices: ( A ) doctor ( B ) bookstore ( C ) market ( D ) train station ( E ) mortuary Only write the answer. Write the answer in the following format: \"Answer: <your answer>\". You must always give an answer. You may only pick one answer choice, if you think multiple are correct only pick the one you think is best.<|eot_id|><|start_header_id|>assistant<|end_header_id|>
CommonsenseQA zero-shot direct answer response for Llama 3.1 70B B